### v1

# CHIPS: Efficient CLIP Adaptation via Curvature-aware Hybrid Influence-based Data Selection

Abstract
Adapting CLIP to vertical domains is typically approached by novel fine-tuning strategies or by continual pre-training (CPT) on large domain-specific datasets. Yet, data itself remains an underexplored factor in this process. We revisit this task from a data-centric perspective: Can effective data selection substitute for large-scale datasets in CPT? We introduce CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), which assigns each image–text pair a utility score that integrates three complementary factors aligned with three goals: faithfulness via a curvature-aware, Newton-style alignment computed in CLIP’s end-point subspace; scalability via an InfoNCE-aware curvature estimator with Johnson–Lindenstrauss (JL) sketching; and retention via a selection-aware relevance weight combined with learnability to balance target adaptation against general-domain preservation. We justify this design theoretically by proving a lower‑bound guarantee on the proxy’s correlation with full‑parameter alignment and by characterizing the bias–variance trade‑offs introduced by curvature mixing and JL sketching. We evaluate CHIPS empirically across various settings: 1) CHIPS attains state-of-the-art performance among selection baselines on 17 medical benchmarks, matches full-dataset CPT with 30% of the data, and outperforms half-dataset CPT using only 10%; 2) on 31 general-domain benchmarks, CHIPS yields the smallest performance drop under 10–30% data-retention budgets. Code, data, and checkpoints will be released.
1 Introduction
Method Target Sketch Curvature Domain CLIP
Full CPT
Random Selection
Rule-filter
CLIPscore
Dot
TracIn
TRAK
CHIPS (ours)
Table 1: Properties of data selection methods for CLIP adaptation. Target: whether the strategy utilizes target evaluation set information. Sketch: whether JL Sketch is supported. Curvature: whether second-order information is employed. Domain: whether general-target domain balance is supported. CLIP: whether the method is specifically designed for CLIP models.
Vision–language models such as CLIP [44] achieve strong zero-shot recognition in general domains, but their performance degrades sharply in vertical settings (e.g., medical imaging and biology) where vocabulary, acquisition protocols, and label taxonomies shift substantially [66]. To adapt CLIP to such domains, two main paradigms have emerged: 1) Model-centric methods modify training or parameterization (e.g., probabilistic fine-tuning [25], many-to-many contrastive learning [24], and PEFT variants [18, 46, 65, 58, 35]); 2) Data-centric methods pursue continual pre-training (CPT) on large, domain-specific datasets, ranging from millions to hundreds of millions of pairs across medical and biodiversity settings [34, 64, 36, 23, 48, 59, 50, 17, 60, 15, 54, 4]. However, for vertical domains, collecting, curating, and processing ever-larger datasets is costly, and indiscriminate upsampling can even dilute learning with redundant, low-utility samples. This naturally raises a question: Is sheer scale truly necessary for effective CPT?
Meanwhile, a complementary line of work studies data attribution [9] to quantify how individual samples affect training dynamics and downstream generalization. Classical influence functions [28] and variants such as TracIn [43], EL2N [42], Arnoldi/GEX/FVM [47, 27, 62], were developed for single-tower model on supervised classification with additive cross-entropy and full-parameter updates. In CLIP, three mechanisms break these assumptions and systematically mis-rank examples. (A) Cross-modal curvature of dual encoders. CLIP’s dual encoders induce non-block-diagonal second-order curvature and block-diagonal proxies ignore this coupling and mis-rank samples. (B) Non-local gradients under InfoNCE. Each sample’s gradient depends on the softmax normalizer over the full negative set, making influence batch-/global-dependent rather than per-example additive and thus sensitive to batch size, queue design, and negative composition. (C) Endpoint dominance at the projection heads. Empirically, the projection heads and temperature drive early shifts in similarity distributions, whereas backbone updates are smaller and slower, rendering full-parameter updates partially unnecessary at the early stage.
Motivated by (A)–(C), we propose CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), a CLIP-specific selector that scores each candidate by the expected one-step drop of the target evaluation loss and selects the top examples for CPT. CHIPS is designed for three goals, faithfulness, scalability, and retention, with one component per goal: (i) a Newton-style alignment computed on the projection heads and temperature that provably lower-bounds full-parameter alignment and enjoys better conditioning (Sec. 2.2); (ii) an InfoNCE-aware curvature preconditioner built from positive/negative gradient moments and compressed via Johnson–Lindenstrauss (JL) sketch to achieve near-linear time/memory with a quantified variance–bias trade-off (Sec. 2.3); and (iii) a selection-aware soft weighting that combines per-sample learnability with target-domain relevance to explicitly control the adaptation–retention balance (Sec. 2.4). Tab. 1 summarizes the dimensions that matter in practice and positions CHIPS against several common baselines.
Across 17 medical benchmarks, CHIPS matches full-dataset CPT using only 30% of the training pool, and outperforms half-dataset CPT using merely 10% data, achieving state-of-the-art among selection baselines. For 31 general-domain benchmarks, although domain adaptation inevitably reduces average performance, CHIPS consistently reduces the drop relative to the previous SOTA selector, indicating stronger retention of general-domain capabilities. In short, with principled attribution and CLIP-aware curvature, effective CPT does not require extreme scale.
Our contributions are threefold:
  • Curvature-aware proxy alignment in CLIP’s end-point subspace. We define a Newton-inspired alignment that ranks samples by expected one-step decrease on the target validation loss and provably lower-bounds full-parameter alignment.
  • InfoNCE-aware curvature estimation with JL sketching and guarantees. We incorporate negative-pair geometry to capture cross-example curvature and characterize the variance–bias trade-off of the estimator.
  • Selection-aware weighting for learnability and retention. We combine learnability and domain-relevance weights to emphasize decision-boundary, on-domain samples while softly limiting selection drift, mitigating catastrophic forgetting of general-domain knowledge.
Figure 1: Workflow of CHIPS. For each training sample, CHIPS computes a curvature-aware proxy Newton alignment in CLIP’s end-point space (projection heads and temperature), where curvature is approximated by mixing self and negative-pair cross moments from symmetric InfoNCE and scaled efficiently via JL sketching. The alignment is then modulated by learnability and target-domain relevance to yield a single selection utility, and the top-$n$ samples are chosen to CPT CLIP models for domain adaptation.
2 Method
CHIPS ranks training samples for CPT by estimating, for each sample, how much a one-step update would decrease the evaluation loss on the target domain. We operationalize this idea along three tightly coupled components: (i) a curvature-aware proxy alignment score computed in CLIP’s end-point geometry (projection heads and temperature), (ii) a negative-pair curvature estimator that captures the second-order coupling induced by symmetric InfoNCE and scales via JL sketching, and (iii) two multiplicative weights for learnability and target-domain relevance, forming a single selection utility. Fig. 1 provides the workflow of CHIPS.
2.1 Problem Setup
We consider a pre-trained general CLIP model with parameters $\theta$ and a large target-domain training pool $\mathcal{D}_{\text{pool}}=\{z_{i}\}_{i=1}^{N}$ of image–text pairs. Our goal is to select a valuable subset from $\mathcal{D}_{\text{pool}}$ for CPT so as to maximize performance on a held-out target-domain evaluation set $\mathcal{D}_{\text{eval}}$ consisting of samples from various validation sets of downstream tasks. Let $\mathcal{L}_{\text{eval}}(\theta)\triangleq\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\ell(z;\theta)]$ denote the evaluation loss, where $\ell$ is the symmetric InfoNCE objective [44] with a learnable logit-scale (temperature) $\tau>0$. Denote the evaluation mean gradient by $\mathbf{u}\triangleq\nabla_{\theta}\mathcal{L}_{\text{eval}}(\theta)=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\nabla_{\theta}\ell(z;\theta)]$. During CPT, one Stochastic Gradient Descent (SGD) step draws a mini-batch of size $B$ from a selection distribution $q$ and computes the batch gradient $\widehat{\mathbf{g}}=\tfrac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\ell(z_{i};\theta)$ to update parameters via
$\theta^{\prime}=\theta-\eta\,\widehat{\mathbf{g}}$, (1)
where $\eta>0$ is the learning rate. We also define the training mean gradient under $q$ by $\mathbf{g}_{q}\triangleq\mathbb{E}_{z\sim q}[\nabla_{\theta}\ell(z;\theta)]$. Assume $\mathcal{L}_{\text{eval}}$ is twice continuously differentiable in a neighborhood of $\theta$ and let $\mathbf{H}_{\text{eval}}(\theta)=\nabla_{\theta}^{2}\mathcal{L}_{\text{eval}}(\theta)$. A second-order Taylor expansion with integral remainder and Hessian-Lipschitz constant $L_{H}$ yields, for $\Delta\theta=-\eta\,\widehat{\mathbf{g}}$,
$\mathbb{E}\big[\Delta\mathcal{L}_{\text{eval}}\big]\;\leq$ $-\eta\,\mathbf{g}_{q}^{\top}\mathbf{u}+\tfrac{1}{2}\,\eta^{2}\,\mathbb{E}\!\big[\,\widehat{\mathbf{g}}^{\!\top}\mathbf{H}_{\text{eval}}(\Theta)\,\widehat{\mathbf{g}}\,\big$] (2)
$+\tfrac{L_{H}}{6}\,\eta^{3}\,\mathbb{E}\!\big[\|\widehat{\mathbf{g}}\|^{3}\big],$
for some $\Theta$ on the segment between $\theta$ and $\theta^{\prime}$. In a local quadratic model, the steepest descent direction is the Newton direction $\mathbf{H}_{\text{eval}}^{-1}\mathbf{u}$, and a good selection distribution $q$ should increase the first-order alignment term $\eta\,\mathbf{g}_{q}^{\top}\mathbf{u}$ while controlling curvature. For a more lightweight assumption, if $\mathcal{L}_{\text{eval}}$ is $\rho$-smooth (gradient-Lipschitz), then
$\mathbb{E}\!\left[\Delta\mathcal{L}_{\text{eval}}\right]\;\leq\;-\eta\,\mathbf{g}_{q}^{\top}\mathbf{u}+\tfrac{\rho}{2}\,\eta^{2}\,\mathbb{E}\!\big[\|\widehat{\mathbf{g}}\|^{2}\big].$ (3)
Eq. (2) and Eq. (3) emphasize the same principle: descent is driven by the alignment between the expected training gradient and the evaluation mean gradient. While we present SGD for clarity, the alignment direction naturally extends to AdamW and full proofs are provided in App. B.1. These observations motivate us to prioritize samples whose updates align with (a proxy of) the Newton direction; the construction of such a curvature-aware proxy in CLIP’s end-point geometry is given next in Sec. 2.2.
2.2 Curvature-aware Proxy Alignment
Given CLIP’s dual-encoder with projection heads and a temperature, we implement CHIPS in the end-point subspace consisting of projection heads and temperature $\vartheta=\{\mathbf{W}_{v},\mathbf{W}_{t},\tau\}$. For a sample $z$, let $\mathbf{g}_{\vartheta}(z)=\nabla_{\vartheta}\ell(z;\theta)$ be its gradient restricted to $\vartheta$, and define the evaluation mean $\mathbf{u}_{\vartheta}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\vartheta}(z)]$. For reference, denote the full-parameter gradient by $\mathbf{g}_{\theta}(z)=\nabla_{\theta}\ell(z;\theta)$ and its evaluation mean by $\mathbf{u}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\theta}(z)]$. Motivated by the one-step descent view in Sec. 2.1, an ideal update direction on $\vartheta$ is the Newton direction $\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$, where $\mathbf{H}_{\vartheta}$ denotes a population curvature (e.g., generalized Gauss–Newton) on the subspace. We therefore define a curvature-aware proxy Newton alignment for each sample:
[…]
where $\mathbf{M}\succ 0$ is a tractable curvature surrogate instantiated in Sec. 2.3. The sign and magnitude of $A(z)$ indicate whether (and how strongly) a one-step update on $z$ moves the model along a descent direction for the evaluation loss: large positive $A(z)$ is preferred.
Directly computing alignment with respect to all parameters is expensive and brittle to high-dimensional noise. We instead argue that the end-point geometry is a reliable proxy for full alignment. Eq. (4) can be rewritten as $A(z)=\tilde{\mathbf{g}}_{\vartheta}(z)^{\top}\tilde{\mathbf{u}}_{\vartheta}$ with $\tilde{\mathbf{g}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{g}_{\vartheta}$ and $\tilde{\mathbf{u}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{u}_{\vartheta}$. Thus $\mathbf{M}$ is a fixed (sample-independent) preconditioner and the proxy–full correlation bound applies unchanged after the congruent transform, i.e., replace $\Sigma_{g}$ by $\mathbf{M}^{-1/2}\Sigma_{g}\mathbf{M}^{-1/2}$ and $\mathbf{S}$ by $\mathbf{M}^{1/2}\mathbf{S}\,\mathbf{M}^{1/2}$. We can state a single unified bound below.
Consider the following local linearization:
$\nabla_{\theta}\ell(z)=\mathbf{J}\,\mathbf{g}_{\vartheta}(z)+\mathbf{r}(z),\qquad\mathbf{u}=\bar{\mathbf{J}}\,\mathbf{u}_{\vartheta}+\boldsymbol{\varepsilon},$ (5)
and define $X(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{u}_{\vartheta}$, $Y(z)=\mathbf{g}_{\theta}(z)^{\top}\mathbf{u}$ with Pearson correlation $\rho_{XY}$, $\Sigma_{g}=\operatorname{Cov}[\mathbf{g}_{\vartheta}(z)]$, $\mathbf{S}=\tfrac{1}{2}(\mathbf{J}^{\top}\bar{\mathbf{J}}+\bar{\mathbf{J}}^{\top}\mathbf{J})$, $\mathbf{B}=\Sigma_{g}^{1/2}\mathbf{S}\,\Sigma_{g}^{-1/2}$, and $\sigma_{\zeta}^{2}=\operatorname{Var}[\zeta(z)]$ for the mismatch term induced by $(\mathbf{r},\boldsymbol{\varepsilon})$.
Theorem 1 (Proxy–full alignment correlation).
If $\zeta(z)$ is uncorrelated with $\mathbf{g}_{\vartheta}(z)$, then
$\rho_{XY}\ \geq\ \frac{\lambda_{\min}\!\big(\mathrm{sym}(\mathbf{B})\big)}{\sqrt{\ \|\mathbf{B}\|_{2}^{2}\;+\;\sigma_{\zeta}^{2}\,/\,\big(\mathbf{u}_{\vartheta}^{\top}\Sigma_{g}\,\mathbf{u}_{\vartheta}\big)\ }}\,.$ (6)
The matrices $\mathbf{S}$ and $\Sigma_{g}$ summarize how end-point gradients relate to full-parameter gradients and how they vary. When $\mathbf{S}$ is well-conditioned along $\mathrm{span}(\mathbf{u}_{\vartheta})$ and the mismatch variance $\sigma_{\zeta}^{2}$ is modest, the bound keeps $\rho_{XY}$ bounded away from zero, so ranking by the proxy alignment tracks the full alignment. Under approximate isotropy or commuting structure, the expression simplifies by effectively replacing $\mathbf{B}$ with $\mathbf{S}$ (see App. B.2). Empirically, on a 100K subset of BIOMEDICA [36], Spearman’s $\rho_{s}=0.83$ between proxy and full alignment supports this proxy view.
2.3 Curvature Estimation
Symmetric InfoNCE couples each positive pair with many negatives through the softmax normalizer, creating cross-example curvature that a positive-only or diagonal surrogate fails to capture. Since our alignment score $A(z)$ relies on a curvature-preconditioned inner product, missing this cross mass yields a biased Newton proxy and unstable rankings along highly coupled directions. Working in the end-point subspace $\vartheta$, we approximate the population curvature (e.g., generalized Gauss–Newton) by combining self and cross second moments of subspace gradients:
$\Phi_{\mathrm{pos}}$ $=\tfrac{1}{N}\sum_{i=1}^{N}\mathbf{g}_{\vartheta}(z_{i})\,\mathbf{g}_{\vartheta}(z_{i})^{\top},$ (7)
$\Phi_{\mathrm{neg}}$ $=\tfrac{1}{N(N-1)}\!\!\sum_{i\neq j}\mathbf{g}_{\vartheta}(z_{i})\,\mathbf{g}_{\vartheta}(z_{j})^{\top},$
where $\Phi_{\mathrm{pos}}$ captures self curvature, while $\Phi_{\mathrm{neg}}$ restores the cross mass induced by random negatives in InfoNCE. We then form a trace- and spectrum-controlled surrogate
$\mathbf{H}_{\vartheta}^{(\alpha)}=(1-\alpha)\,\Phi_{\mathrm{pos}}+\alpha\,\Phi_{\mathrm{neg}},\qquad\mathbf{M}=\mathbf{H}_{\vartheta}^{(\alpha)}+\lambda\,\mathbf{I},$ (8)
with a mixing weight $\alpha\in[0,1]$ and a small Tikhonov parameter $\lambda>0$ to ensure numerical stability in matrix inversion. Plugging $\mathbf{M}$ into Eq. (4) yields the curvature-aware proxy Newton score $A(z)$. By mixing $\Phi_{\mathrm{pos}}$ with the cross moment $\Phi_{\mathrm{neg}}$ (with $\alpha>0$), we restore the off-diagonal curvature induced by InfoNCE’s negatives that positive-only surrogates ignore. The surrogate $\mathbf{M}\!\succ\!0$ thus preconditions $A(z)$ in a fixed metric without altering the proxy–full correlation framework (Sec. 2.2); $\alpha$ controls coupling strength and $\lambda$ provides ridge stability. This yields an alignment score that is sensitive to the true second-order structure yet robust. Let the ideal scalar be $A^{\star}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$ and the curvature-approximated score $A_{\alpha}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$ with $\mathbf{M}$ as in Eq. (8), and define $\Delta_{\alpha}=\mathbf{H}_{\vartheta}-\mathbf{H}_{\vartheta}^{(\alpha)}$. For further efficiency, we apply a $k$-dimensional JL sketch to compute a sketched score $\widehat{A}_{\alpha}(z)$. We discuss the choice of JL sketch in Sec. 4.1.
Theorem 2 (Error bound for curvature mixing).
Assume bounded spectra for $\mathbf{H}_{\vartheta}$ and $\mathbf{H}_{\vartheta}^{(\alpha)}$. Then, with probability at least $1-\delta$ over the JL sketch (vacuous if no sketch is used), there exist constants $C_{1},C_{2}>0$, depending only on spectral bounds of $\mathbf{H}_{\vartheta}$ and $\mathbf{H}_{\vartheta}^{(\alpha)}$, such that
$\mathbb{E}_{z}\!\big[\big(\widehat{A}_{\alpha}(z)-A^{\star}(z)\big)^{2}\big]$ $\leq\underbrace{\frac{C_{1}\log(1/\delta)}{k}}_{\text{projection variance}}$ (9)
$\quad+\underbrace{C_{2}\,\|\Delta_{\alpha}\|_{F}^{2}\,\|\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}\|_{2}^{2}}_{\text{curvature bias}}.$
Eq. (9) separates an $O(1/k)$ projection variance term and a curvature bias term that shrinks as $\mathbf{H}_{\vartheta}^{(\alpha)}$ better matches $\mathbf{H}_{\vartheta}$. Because $\Phi_{\mathrm{neg}}$ restores off-diagonal mass whenever cross-example coupling is present, any $\alpha>0$ that does so reduces $\|\Delta_{\alpha}\|_{F}$ and tightens the bound; $\lambda$ stabilizes the inverse without changing the decomposition. All derivations and proofs are deferred to App. B.3.
2.4 Learnability and Domain Relevance
The curvature-aware alignment score $A(z)$ in Eq. (4) indicates whether a sample points in a useful descent direction, but it does not distinguish already-solved cases from borderline ones, nor does it compensate for the distributional gap between $\mathcal{D}_{\text{pool}}$ and $\mathcal{D}_{\text{eval}}$. Therefore, we introduce another two multiplicative weights to modulate $A(z)$ by learnability and target-domain relevance.
Learnability.
For a mini-batch of paired image–text examples $\{(x_{i},y_{i})\}_{i=1}^{B}$, let the end-point embeddings be $\hat{\mathbf{x}}_{i}=\frac{\mathbf{W}_{v}\mathbf{h}_{i}}{\|\mathbf{W}_{v}\mathbf{h}_{i}\|}$ and $\hat{\mathbf{y}}_{j}=\frac{\mathbf{W}_{t}\mathbf{t}_{j}}{\|\mathbf{W}_{t}\mathbf{t}_{j}\|}$ with temperature $\tau>0$. The CLIP similarity is $s_{ij}=\tau\,\hat{\mathbf{x}}_{i}^{\top}\hat{\mathbf{y}}_{j}$, with softmax probabilities $p^{i2t}_{ij}=\frac{\exp(s_{ij})}{\sum_{j^{\prime}}\exp(s_{ij^{\prime}})}$ and $p^{t2i}_{ij}=\frac{\exp(s_{ij})}{\sum_{i^{\prime}}\exp(s_{i^{\prime}j})}$. Define the average correctness and a hardest-negative margin as
$p_{\mathrm{corr}}(z)$ $=\tfrac{1}{2}\big(p^{i2t}_{ii}+p^{t2i}_{ii}\big),$ (10)
$m(z)$ $=s_{ii}-\max\!\big\{\max_{j\neq i}s_{ij},\,\max_{i^{\prime}\neq i}s_{i^{\prime}i}\big\}.$
We then weight toward decision-boundary examples and away from saturated ones:
$w_{\mathrm{L}}(z)=\big(1-p_{\mathrm{corr}}(z)\big)\,\big(1+\sigma(-m(z))\big),$ (11)
where $\sigma(\cdot)$ is the sigmoid function. The first factor downweights already-correct and high-confidence pairs to reduce their impact while the second emphasizes small/negative margins that are most learnable in one step.
Target-domain Relevance.
To softly favor samples consistent with $\mathcal{D}_{\text{eval}}$, we compute evaluation embeddings $\boldsymbol{\mu}_{x}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\hat{\mathbf{x}}$], $\boldsymbol{\mu}_{y}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\hat{\mathbf{y}}$], and define
$w_{\mathrm{R}}(z)=\sigma\!\Big((1-\beta)\,\cos(\hat{\mathbf{x}},\boldsymbol{\mu}_{x})+\beta\,\cos(\hat{\mathbf{y}},\boldsymbol{\mu}_{y})\Big),$ (12)
where $\beta\in[0,1]$ controls the balance between two modalities and $\cos(\cdot,\cdot)$ denotes the cosine similarity. Notably, the relevance logit $s(z)=(1-\beta)\,\cos(\hat{\mathbf{x}},\boldsymbol{\mu}_{x})+\beta\,\cos(\hat{\mathbf{y}},\boldsymbol{\mu}_{y})$ in Eq. (12) lies in [1,1][-1,1], therefore $w_{\mathrm{R}}(z)\in[\sigma(-1),\sigma(1)]\approx[0.27,0.73]$. This implements soft re-weighting to some extent rather than hard filtering as no sample is zeroed out.
Let $\tilde{q}\propto A(\cdot)\,w_{\mathrm{L}}(\cdot)$ and $q\propto w_{\mathrm{R}}(\cdot)\,\tilde{q}$ be the selection distributions with/without $w_{\mathrm{R}}$. Because $w_{\mathrm{R}}\in[a,b]$ with $a=\sigma(-1),\,b=\sigma(1)$, the density ratio $r(z)=\frac{q(z)}{\tilde{q}(z)}=\frac{w_{\mathrm{R}}(z)}{\mathbb{E}_{\tilde{q}}[w_{\mathrm{R}}]}$ satisfies $r(z)\in[a/b,\,b/a]=[e^{-1},e]$, which implies a bounded selection drift $D_{\mathrm{KL}}(q\,\|\,\tilde{q})\leq 1$ nat. Thus $w_{\mathrm{R}}$ also limits how far the selection moves from the base alignment-driven distribution, mitigating catastrophic forgetting of general-domain knowledge. We analyze $\beta$ in Sec. 4.2 to balance adaptation and retention.
Final utility score and selection.
We combine the components multiplicatively into a single ranking score.
$\mathcal{I}_{\text{CHIPS}}(z)=\widehat{A}_{\alpha}(z)\cdot w_{\mathrm{L}}(z)\cdot w_{\mathrm{R}}(z),$ (13)
and select the top-$n$ examples from $\mathcal{D}_{\text{pool}}$ for CPT. Implementation details of CHIPS are provided in App. C.
3 Experiment
Model Medical General
OPH RAD DER HEM PAT NEU HIS BIO Avg CLS RET
PubMedCLIP 35.92 20.27 9.63 8.45 25.77 60.50 8.45 25.27 26.21 19.05 24.21
BioMedCLIP 14.42 15.59 1.20 10.14 21.17 44.79 4.66 72.91 21.45 7.68 23.55
BMCLIP 38.47 17.44 61.65 9.44 49.75 38.26 5.02 17.28 29.86 57.97 28.27
Vanilla 15.91 19.76 10.82 17.95 30.09 57.96 5.24 10.56 23.51 53.15 29.23
Full Dataset 33.30 12.75 11.72 23.33 49.57 82.79 12.40 9.81 31.51 49.72 24.20
rr=10%
Random 36.01 9.93 11.27 20.61 39.71 43.96 4.68 10.18 24.78 52.21 29.28
Concept-Balance 33.18 12.11 11.67 21.81 40.40 46.82 5.03 10.87 25.50 51.95 30.14
Concept-Filter 31.40 10.81 11.72 23.41 43.14 44.53 5.14 11.06 25.15 51.81 29.19
CLIPScore 15.48 12.75 11.47 20.99 38.23 64.15 5.25 10.56 24.16 53.39 31.27
Dot 42.01 8.54 12.17 16.87 36.98 45.49 4.16 23.51 25.32 48.27 26.47
TracIn 41.30 10.47 12.02 16.69 35.69 51.79 4.96 26.34 26.46 47.26 26.36
TRAK 30.29 12.37 12.17 25.93 39.04 45.95 4.46 12.88 25.19 48.24 27.17
CHIPS (ours) 38.93 11.05 11.87 21.43 37.99 50.22 5.29 25.96 27.03 47.88 25.71
rr=20%
Random 35.79 13.82 11.62 21.92 39.35 36.41 6.60 11.13 25.00 51.39 29.57
Concept-Balance 32.77 13.62 11.97 24.00 44.24 43.16 5.42 10.37 26.20 51.38 30.31
Concept-Filter 31.36 11.70 11.97 25.46 44.98 35.85 6.42 15.21 25.23 51.25 29.54
CLIPScore 17.33 9.83 11.27 18.82 38.01 32.35 4.45 11.25 20.01 51.25 28.81
Dot 39.20 11.19 11.77 13.30 35.46 53.72 5.38 33.63 26.39 47.50 26.93
TracIn 42.47 12.28 11.82 19.70 37.92 46.13 5.07 19.17 26.63 46.79 25.09
TRAK 43.25 10.12 11.02 16.75 37.66 28.83 5.41 28.16 24.54 47.54 25.81
CHIPS (ours) 44.13 13.02 11.87 22.49 40.79 38.54 5.92 30.36 28.20 47.45 25.90
rr=30%
Random 36.77 11.90 12.07 18.42 43.35 47.08 6.42 10.94 26.28 50.98 29.85
Concept-Balance 31.95 11.88 12.07 24.61 42.52 35.88 6.48 10.43 24.56 51.13 29.66
Concept-Filter 32.37 11.23 12.02 25.96 41.40 41.02 6.81 21.06 25.37 50.63 29.54
CLIPScore 21.93 8.25 11.32 13.94 38.81 26.49 4.49 11.94 19.01 50.76 28.09
Dot 38.42 11.05 11.62 15.40 37.52 35.04 7.69 20.74 23.97 46.39 25.67
TracIn 39.63 14.18 11.97 14.24 38.91 41.09 5.69 22.19 25.68 46.23 25.15
TRAK 38.95 10.49 11.42 16.92 36.96 25.70 7.75 25.14 23.54 46.29 24.04
CHIPS (ours) 41.66 13.56 11.37 25.02 43.10 46.53 7.69 36.34 29.96 46.34 26.11
rr=50%
Random 34.47 13.86 11.82 18.82 45.82 41.00 10.06 10.56 26.26 50.56 29.29
Table 2: Downstream performance of MetaCLIP-B16-400M [57] and variants continually pre-trained with data subsets from BIOMEDICA [36] selected by different methods, under a fixed retention ratio $r\in\{10\%,20\%,30\%,50\%\}$. Abbreviations: for medical specialties, OPH = Ophthalmology, RAD = Radiology, DER = Dermatology, HEM = Hematology, PAT = Pathology, NEU = Neuropathology, HIS = Histology, BIO = Non-clinical Biology; for metrics in general domain, CLS denotes classification accuracy, and RET denotes the average R@1 across image-to-text and text-to-image retrieval tasks. In each setting, the best result is bolded and the second best is underlined.
3.1 Experimental Setup
Training.
For a comprehensive comparison, we perform continual pre-training using the initialization weights of MetaCLIP-B32/B16/L14/H14 [57]. We select medical as target domain and two recently introduced large-scale medical multimodal datasets are employed as training pools: BIOMEDICA (24M samples) [36] and MedTrinity (18M samples) [56]. We select data retention ratio $r$ as 10%, 20%, 30%, and 50% of the complete dataset, with the number of training epochs fixed at 5. All models are optimized using the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=10^{-6}$) and a cosine learning rate scheduler (initial learning rate as $10^{-6}$), with a global batch size of 32,768. Each experiment is conducted on an Ubuntu 24.04 server equipped with 8x NVIDIA H200 (141GB) GPUs. Further details regarding the training datasets, model architectures, and specific hyperparameter settings are provided in App. E.
Evaluation.
We evaluate 48 tasks spanning general and medical domains to enable a comprehensive assessment. The general-domain suite includes Cars [29], Country211 [44], Fer2013 [16], Aircraft [37], Food101 [3], GTSRB [49], ImageNet-A, ImageNet-O [20], ImageNet-1K [8], ImageNetV2 [45], MNIST [31], Rendered-SST2 [44], STL-10 [7], SUN397 [55], VOC [12], Caltech-101 [13], CIFAR-10, CIFAR-100 [30], CLEVR (Closest-Object-Distance/Count-All) [26], DTD [6], EuroSAT [19], Oxford Flowers [39], KITTI [14], Pets [41], RESISC45 [5], smallNORB (Azimuth), smallNORB (Elevation) [32], SVHN [38], Flickr8k [22], Flickr30k [63], and MSCOCO [33]. For the medical domain, we consider 17 classification tasks across eight specialties. These include Ophthalmology (Diabetic [11], OCTMNIST, RetinaMNIST [61]); Radiology (ChestMNIST, ChestX-ray14 [52], OrganAMNIST, OrganCMNIST, OrganSMNIST); Dermatology (DermaMNIST); Hematology (BloodMNIST); Pathology (PCAM [51], LC25000 [2], PathMNIST); Neuropathology (Amyloid CAA/Diffuse [53]); Histology (TissueMNIST); and Non-clinical Biology (Pollen [1]). For classification tasks, the average accuracy is reported. For retrieval tasks, bidirectional (image-to-text and text-to-image) performance is evaluated using $R@1$ recall metric. Further details are provided in App. F.
Baselines.
We select a total of seven data selection baselines for comparison, including (1) random selection, which selects samples uniformly from the training pool without any selective control; (2) three heuristics-based methods: CLIPScore [21], Concept-Balance, and Concept-Filter [36]; (3) Influence-based methods: Dot [67], TracIn [43], and TRAK [40]. Moreover, we also include three commonly used medical-specific CLIP models for reference: PubMedCLIP [10], BioMedCLIP [64], and BMCLIP [36]. Implementation details of all baseline methods and our CHIPS method are provided in App. C.
Model Medical CLS General CLS General RET
B32-400M
Random 27.15 49.31 27.33
TracIn 27.48 47.19 25.10
CHIPS (ours) 27.83 47.90 25.65
B32-$C$
Random 27.95 50.69 29.28
TracIn 26.25 48.90 25.72
CHIPS (ours) 28.13 49.48 26.79
B16-400M
Random 24.78 52.21 29.28
TracIn 26.46 47.26 26.36
CHIPS (ours) 27.03 47.88 25.71
B16-$C$
Random 26.30 55.91 30.16
TracIn 25.89 50.17 27.37
CHIPS (ours) 26.93 51.27 28.17
L14-400M
Random 29.33 57.07 33.35
TracIn 27.08 52.99 28.17
CHIPS (ours) 29.73 53.65 28.17
L14-$C$
Random 30.75 60.74 32.34
TracIn 31.54 56.98 28.26
CHIPS (ours) 31.74 57.93 30.05
H14-$C$
Random 35.23 61.36 32.82
TracIn 35.10 58.27 31.60
CHIPS (ours) 35.48 58.24 32.09
Table 3: Generalization experiment of MetaCLIP-series adapted models continually pre-trained on 10% kept data from different data selection methods. 400M denotes the model was pre-trained on 400M general image-text pairs and $C$ denotes the model was pre-trained on 2.5B general image-text pairs.
3.2 Main Results
Effectiveness of CHIPS.
As shown in Tab. 2, across all retention ratios, CHIPS achieves the best average performance on 17 medical tasks among data selection methods, with Medical Avg scores of 27.03, 28.20, and 29.96, exceeding the second–best method by +0.57, +1.57, and +3.68 points, respectively. Notably, with only 10% of the full BIOMEDICA, CHIPS outperforms a 50% random subset (27.03 vs. 26.26); with 30%, it achieves 95.1% of the full-dataset performance (29.96 vs. 31.51). It is also competitive with specialized medical CLIP models pre-trained on larger datasets: at $r=30\%$, CHIPS slightly surpasses BMCLIP (29.96 vs. 29.86) and consistently outperforms PubMedCLIP and BioMedCLIP across $r$. Even though CHIPS is not uniformly best on every specialty, it yields the highest overall average. Importantly, CHIPS better preserves general-domain abilities than the previous SOTA selection method TracIn: for CLS, CHIPS retains 90.1%, 89.3%, and 87.2% of the vanilla model at $r=10\%,20\%,30\%$ (vs. 88.9%, 88.0%, 87.0% for TracIn), and for RET it is comparable, slightly lower at $r=10\%$ but higher at $r\geq 20\%$, yielding a higher average RET across retention ratios. Overall, CHIPS delivers the strongest medical performance while incurring less degradation on the general domain than TracIn.
Generalization abilities of CHIPS.
With a single scoring pass on MetaCLIP-B16-400M (10% kept), we reuse the computed CHIPS scores to train diverse backbones (B32/B16/L14/H14) and pre-training scales (400M and CC). As shown in Tab. 3, across all seven settings, CHIPS attains the best Medical performance, outperforming TracIn by 0.20-2.65 points. On general-domain metrics, CHIPS typically ranks second (behind Random) and generally matches or exceeds TracIn on CLS and RET, with a single RET exception on B16-400M. Thus, CHIPS transfers across architectures and scales, allowing cached scores to be reused while delivering strong medical gains with minimal loss of general ability.
Ablation study.
As shown in Tab. 4, we ablate CHIPS by progressively adding components from Eq. (13). On the medical benchmarks, CHIPS achieves the highest performance at all budgets, surpassing the strongest ablation by +1.05, +0.28, and +1.46 points at $r=10\%,20\%,30\%$, respectively. These gains support the multiplicative combination of alignment, learnability, and domain relevance, with the relevance term becoming especially beneficial at larger budgets ($r=30\%$). Meanwhile, general-domain performance remains close to the best ablation (within $\leq$0.53 for General CLS and $\leq$0.99 for General RET), indicating minimal trade-off for the medical improvements. On the general domain, the retrieval gap to the best ablation narrows as $r$ grows (0.99 $\rightarrow$ 0.66 $\rightarrow$ 0.37 for General RET), suggesting controlled specialization towards the target domain rather than catastrophic forgetting.
Cost analysis.
We measure scoring cost as total FLOPs over $\mathcal{D}_{\text{pool}}$. CHIPS requires $50.9475{\times}10^{15}$ FLOPs, 3.1% lower than TracIn ($52.5891{\times}10^{15}$) and effectively on par with TRAK ($50.9458{\times}10^{15}$). Despite equal-or-lower cost, CHIPS yields consistently stronger medical performance than TracIn (Tab. 2), improving by +0.57, +1.57, and +4.28 points across three retention settings, respectively. In practice, the one-time scoring overhead is further amortized because CHIPS scores can be cached and reused across architectures and pre-training scales (Tab. 3). The full FLOPs computation process is provided in App. G.
Model Medical CLS General CLS General RET
r=10%
Alignment-only 25.98 48.33 26.70
Alignment-Margin 25.95 48.41 26.25
CHIPS (ours) 27.03 47.88 25.71
r=20%
Alignment-only 27.52 46.67 26.25
Alignment-Margin 27.92 46.88 26.56
CHIPS (ours) 28.20 46.73 25.90
r=30%
Alignment-only 27.84 46.78 25.23
Alignment-Margin 28.50 46.75 25.52
CHIPS (ours) 29.96 46.34 25.15
Table 4: Ablation experiment of CHIPS on MetaCLIP-B16-400M under 10%, 20%, and 30% data retention ratios. Alignment-only uses single alignment score, Alignment-Margin additionally introduces the margin term, and CHIPS multiplies all three metrics.
4 Analysis
4.1 Alignment
Figure 2: Downstream results of MetaCLIP-B16-400M continually pre-trained on 10% data selected from CHIPS under: (A) different evaluation set sizes, (B) different mixing $\alpha$ in computing alignment scores, and (C) different balance $\beta$ in computing relevance scores.
Effect of target evaluation set size.
Expanding the target evaluation set $\mathcal{D}_{\text{eval}}$ sharpens the alignment direction $\mathbf{u}_{\vartheta}$ and improves medical accuracy with minimal general-domain drift. As shown in Fig. 2(A), as samples per task increase from 50 to 200, Medical CLS improves from 26.25 to 27.03 (+0.78), then saturates at 250 (27.12, +0.09). The slight dip at 100 (25.82) reflects higher variance at smaller set sizes, while overall the trend is upward and stabilizing around 200. General CLS remains essentially flat across all sizes (48.41 at 50 to 47.88 at 200; total range $\leq$0.55), indicating that larger $\mathcal{D}_{\text{eval}}$ chiefly benefits medical alignment without harming general performance. Balancing diminishing returns beyond 200 against the linear increase in scoring cost, and we adopt 200 samples per task as the default.
Figure 3: Medical downstream results of MetaCLIP-B16-400M continually pre-trained on 10% data selected from different selection methods under various end-point geometry settings.
Projection layers dominate alignment with text being more informative.
We ablate the end-point geometry $\vartheta=\{\mathbf{W}_{v},\mathbf{W}_{t},\tau\}$ into All, Logit-only, Visual-only, and Text-only, with results shown in Fig. 3. Across methods, Text-only nearly matches All (98.6–99.7%), Visual-only lags, and Logit-only is weakest. For CHIPS, the scores are 27.03 (All), 26.95 (Text-only, -0.08; 99.7%), 26.67 (Visual-only, -0.36; 98.7%), and 23.83 (Logit-only, -3.20). These results show that alignment resides primarily in the projection heads, especially $\mathbf{W}_{t}$, with $\mathbf{W}_{v}$ providing a smaller yet complementary gain. Moreover, CHIPS is more robust to projection layers compared with baseline methods, retaining 99.7% (Text-only) and 98.7% (Visual-only), the highest among methods, and exhibiting the smallest text–visual gap (0.28 vs. 2.43/3.77/1.05 for Dot/TracIn/TRAK).
Effect of random projection.
[…]
Balancing positive and negative curvature.
In Eq. (8), $\alpha$ linearly mixes positive and negative pair curvature to form $\mathbf{M}$, which reweights gradients in $A(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$. As shown in Fig. 2(B), Medical CLS peaks at $\alpha{=}0.6$ (27.05) and remains essentially tied at $0.8$ (27.03), while both smaller ($0.2$: 25.92; $0.4$: 25.66) and larger ($1.0$: 26.07) values underperform by about 1–1.4 points. General CLS is almost flat across $\alpha$ (48.36–48.46), with a modest dip at $0.6$ (47.88), indicating that $\alpha$ mainly governs target-domain discrimination with negligible general-domain impact. Mechanistically, moderate $\alpha$ balances attraction to positives and repulsion from negatives, improving the conditioning of $\mathbf{M}$ and yielding a more discriminative alignment direction; extremes overweight one term and reduce $A(z)$’s utility. We therefore recommend $\alpha\in[0.6,0.8]$ (default $0.6$ for medical focus; $0.4$ if slightly prioritizing general-domain retention).
Figure 4: Medical downstream results of MetaCLIP-B16-400M continually pre-trained on 10% data selected from CHIPS under various random projection settings.
4.2 Domain Relevance and General Retention
[…]
5 Conclusion
We propose CHIPS, showing that principled data selection can substitute for raw scale in CLIP continual pre-training. On 17 medical benchmarks, CHIPS attains SOTA performance among selection baselines, matches full-dataset CPT with 30% of the data, and outperforms half-dataset CPT using only 10%. Across 31 general-domain benchmarks, it yields the smallest performance drop under 10–30% retention ratios, better preserving general capabilities. Limitations include reliance on a target validation distribution, we will explore unlabeled or shift-robust target signals and extensions beyond the CLIP model in the future work.
Supplementary Material
Appendix Contents
Appendix A Related Work
CLIP Adaptation.
Current approaches for effective CLIP adaptation to specialized vertical domains (e.g., Medical) can be broadly categorized into two paradigms: model-centric and data-centric methods [66]. Model-centric approaches primarily focus on developing novel training strategies, including probabilistic fine-tuning [25] and many-to-many contrastive learning [24], as well as parameter-efficient fine-tuning (PEFT) techniques [18, 46, 65, 58]. In contrast, data-centric approaches emphasize the collection of large-scale domain-specific datasets for continual pre-training. Within the general-purpose medical domain, PMC-CLIP [34] leveraged approximately 2M samples, BioMedCLIP [64] curated 15M samples, BIOMEDICA [36] assembled 24M samples, and MedTrinity aggregated 25M samples. For more specialized medical subdomains, QUILT [23] compiled 1M samples for Histopathology, EYECLIP [48] gathered approximately 3M samples for Ophthalmology, and PanDerm [59] curated 2M samples for Dermatology. In the biology and biodiversity domains, even larger-scale datasets ranging from 5M to 214M samples have been collected [50, 17, 60, 15, 54, 4]. In this work, rather than pursuing resource-intensive dataset scaling, we investigate the intrinsic characteristics of existing domain-specific datasets and employ strategic data selection to achieve more efficient CLIP adaptation.
Data Attribution.
Data attribution methods quantify the influence of individual training samples on model training efficacy. [28] initially introduced Influence Functions (IF), proposing the LiSSA approximation to efficiently compute the inverse Hessian-vector product, wherein the computation entails calculating the Hessian inverse, iterating through validation and training samples, and deriving influence scores through gradient-based calculations. [43] proposed TracIn, which offers a more intuitive first-order approximation that tracks model training dynamics across multiple checkpoints. [42] introduced EL2N, a self-influence metric that computes the L2 norm of prediction error vectors without requiring explicit validation samples. [47] proposed the Arnoldi iteration method, which leverages dominant eigenvalue decomposition to address the computational overhead and non-H-invariance issues inherent in random projection approaches. [27] introduced GEX, reformulating IF to capture bimodal influence distributions by replacing gradient-based Taylor approximations with direct loss multiplications and employing ensemble models to mitigate Hessian singularity bias. [62] proposed FVM, which optimizes for flat validation minima to enhance the stability of influence estimations for mislabeled sample detection. In this work, we propose an IF variant specifically tailored for CLIP influence estimation.
Appendix B Proofs
This section provides consolidated derivations and proofs for the three core components referenced in Sec. 2.1, Sec. 2.2, and Sec. 2.3. Unless stated otherwise, expectations are taken over the mini-batch sampling and data randomness at the current parameter iterate $\theta$. We write $\widehat{\mathbf{g}}=\tfrac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\ell(z_{i};\theta)$, $\mathbf{g}_{q}=\mathbb{E}_{z\sim q}[\nabla_{\theta}\ell(z;\theta)]$, $\mathbf{u}=\nabla_{\theta}\mathcal{L}_{\text{eval}}(\theta)$, and use the Euclidean inner product and norm. Spectral and Frobenius norms are denoted by $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$.
B.1 Alignment Direction
In a neighborhood of $\theta$ assume either (i) $\mathcal{L}_{\text{eval}}\in C^{2}$ with Hessian-Lipschitz constant $L_{H}$, or (ii) $\mathcal{L}_{\text{eval}}$ is $\rho$-smooth.
Second-order upper bound.
For $\Delta\theta$, the second-order Taylor expansion with integral remainder gives
$\mathcal{L}_{\text{eval}}(\theta+\Delta\theta)$ $=\mathcal{L}_{\text{eval}}(\theta)+\mathbf{u}^{\top}\Delta\theta$
$\quad+\tfrac{1}{2}\,\Delta\theta^{\top}\mathbf{H}_{\text{eval}}(\Theta)\,\Delta\theta+R_{3},$ (14)
where $\Theta$ lies on the segment between $\theta$ and $\theta+\Delta\theta$, and $|R_{3}|\leq\tfrac{L_{H}}{6}\,\|\Delta\theta\|^{3}$. With the one-step SGD move $\Delta\theta=-\eta\,\widehat{\mathbf{g}}$ and taking expectation,
$\mathbb{E}\big[\Delta\mathcal{L}_{\text{eval}}\big]$ $\leq-\eta\,\mathbf{g}_{q}^{\top}\mathbf{u}$
$\quad+\tfrac{1}{2}\,\eta^{2}\,\mathbb{E}\!\big[\widehat{\mathbf{g}}^{\!\top}\mathbf{H}_{\text{eval}}(\Theta)\widehat{\mathbf{g}}\big$]
$\quad+\tfrac{L_{H}}{6}\,\eta^{3}\,\mathbb{E}\!\big[\|\widehat{\mathbf{g}}\|^{3}\big],$
which matches Eq. (2) in the main text and shows that descent is driven by the alignment term $\mathbf{g}_{q}^{\top}\mathbf{u}$.
First-order upper bound.
If $\mathcal{L}_{\text{eval}}$ is $\rho$-smooth, the Descent Lemma yields
$\mathcal{L}_{\text{eval}}(\theta+\Delta\theta)\leq\mathcal{L}_{\text{eval}}(\theta)+\mathbf{u}^{\top}\Delta\theta+\tfrac{\rho}{2}\|\Delta\theta\|^{2}.$ (15)
Setting $\Delta\theta=-\eta\,\widehat{\mathbf{g}}$ and taking expectations gives Eq. (3).
Reegarding mini-batch variance, let $\Sigma_{q}=\operatorname{Cov}_{z\sim q}[\nabla_{\theta}\ell(z;\theta)]$, then
$\mathbb{E}\big[\|\widehat{\mathbf{g}}\|^{2}\big]$ $=\|\mathbf{g}_{q}\|^{2}+\tfrac{1}{B}\operatorname{tr}(\Sigma_{q}),$ (16)
$\mathbb{E}\big[\|\widehat{\mathbf{g}}-\mathbf{g}_{q}\|^{2}\big]$ $=\tfrac{1}{B}\operatorname{tr}(\Sigma_{q}).$
This makes explicit how the batch size $B$ moderates the quadratic term in Eq. (3). In a local quadratic model the steepest descent direction for $\mathcal{L}_{\text{eval}}$ is $\mathbf{H}_{\text{eval}}^{-1}\mathbf{u}$. Hence a selection distribution $q$ that increases $\mathbf{g}_{q}^{\top}\mathbf{H}_{\text{eval}}^{-1}\mathbf{u}$ is desirable, which motivates the proxy alignment in Sec. 2.2.
AdamW-aware form.
At iteration $t$, let $\mathbf{g}_{t}=\tfrac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\ell(z_{i};\theta_{t})$. Consider Adam moments and bias corrections as the following forms
$m_{t}$ $=\beta_{1}m_{t-1}+(1-\beta_{1})\mathbf{g}_{t},$
$v_{t}$ $=\beta_{2}v_{t-1}+(1-\beta_{2})\,(\mathbf{g}_{t}\odot\mathbf{g}_{t}),$
$\hat{m}_{t}$ $=\frac{m_{t}}{1-\beta_{1}^{t}},$
$\hat{v}_{t}$ $=\frac{v_{t}}{1-\beta_{2}^{t}},$
$\mathbf{P}_{t}$ $=\operatorname{diag}\big((\sqrt{\hat{v}_{t}}+\epsilon)^{-1}\big),$
where $\odot$ denotes the element-wise product. With decoupled weight decay $w_{d}>0$ and diagonal mask $\mathbf{D}$, the update is
$\Delta\theta_{t}\triangleq\theta_{t+1}-\theta_{t}=-\eta_{t}\big(\mathbf{P}_{t}\hat{m}_{t}+w_{d}\,\mathbf{D}\,\theta_{t}\big).$
Under the $C^{2}$ and Hessian-Lipschitz assumption, Eq. (B.1) gives
$\mathbb{E}\big[\Delta\mathcal{L}_{\text{eval}}\big]\leq\,$ $-\eta_{t}\,\mathbb{E}\big[\hat{m}_{t}^{\top}\mathbf{P}_{t}\mathbf{u}\big]-\eta_{t}\,w_{d}\,(\mathbf{D}\theta_{t})^{\top}\mathbf{u}$
$+\tfrac{1}{2}\,\eta_{t}^{2}\,\mathbb{E}\!\big[\Delta\theta_{t}^{\top}\mathbf{H}_{\text{eval}}(\Theta_{t})\Delta\theta_{t}\big]$
$+\tfrac{L_{H}}{6}\,\eta_{t}^{3}\,\mathbb{E}\!\big[\|\Delta\theta_{t}\|^{3}\big],$
Moreover, if under $\rho$-smoothness, Eq. (15) yields
$\mathbb{E}\big[\Delta\mathcal{L}_{\text{eval}}\big]$ $\leq-\eta_{t}\,\mathbb{E}\big[(\mathbf{P}_{t}\hat{m}_{t})^{\top}\mathbf{u}\big]$
$\quad-\eta_{t}\,w_{d}\,(\mathbf{D}\theta_{t})^{\top}\mathbf{u}$
$\quad+\tfrac{\rho}{2}\,\eta_{t}^{2}\,\mathbb{E}\!\big[\|\Delta\theta_{t}\|^{2}\big].$
In both cases the alignment term becomes $\mathbb{E}[(\mathbf{P}_{t}\hat{m}_{t})^{\top}\mathbf{u}]$, which is in line with SDG in the main content of this paper.
B.2 Proxy Alignment in Subspace
In this section, we prove Theorem 1 and state two practical corollaries.
In the end-point subspace let $\mathbf{g}=\mathbf{g}_{\vartheta}(z)$, $\boldsymbol{\mu}=\mathbb{E}[\mathbf{g}]$, $\mathbf{g}_{c}=\mathbf{g}-\boldsymbol{\mu}$, and $\Sigma_{g}=\operatorname{Cov}[\mathbf{g}]\succeq 0$. Locally,
$\nabla_{\theta}\ell(z)=\mathbf{J}\,\mathbf{g}+\mathbf{r}(z),\hskip 28.80008pt\mathbf{u}=\bar{\mathbf{J}}\,\mathbf{u}_{\vartheta}+\boldsymbol{\varepsilon},$
with $\mathbf{J},\bar{\mathbf{J}}$ linear maps, residual $\mathbf{r}(z)$, and deterministic mismatch $\boldsymbol{\varepsilon}$. Define
$\mathbf{S}=\tfrac{1}{2}(\mathbf{J}^{\top}\bar{\mathbf{J}}+\bar{\mathbf{J}}^{\top}\mathbf{J}),\qquad\mathbf{A}=\tfrac{1}{2}(\mathbf{J}^{\top}\bar{\mathbf{J}}-\bar{\mathbf{J}}^{\top}\mathbf{J}),$
and consider
$X(z)=\mathbf{u}_{\vartheta}^{\top}\mathbf{g}_{c},\hskip 28.80008ptY(z)=(\nabla_{\theta}\ell(z))^{\top}\mathbf{u}.$
Collect the remaining random terms into
$\zeta(z)=(\mathbf{J}\mathbf{g}_{c})^{\top}\boldsymbol{\varepsilon}+\mathbf{r}(z)^{\top}\bar{\mathbf{J}}\mathbf{u}_{\vartheta}+\mathbf{r}(z)^{\top}\boldsymbol{\varepsilon}+\mathbf{g}_{c}^{\top}\mathbf{A}\,\mathbf{u}_{\vartheta},$
and write $\sigma_{\zeta}^{2}=\operatorname{Var}[\zeta(z)]$.
Proof of Theorem 1.
Let $\tilde{\mathbf{g}}=\Sigma_{g}^{-1/2}\mathbf{g}_{c}$, $\alpha=\Sigma_{g}^{1/2}\mathbf{u}_{\vartheta}$, $B=\Sigma_{g}^{1/2}\mathbf{S}\Sigma_{g}^{-1/2}$, and $Z(z)=\mathbf{g}_{c}^{\top}\mathbf{S}\mathbf{u}_{\vartheta}=\alpha^{\top}B\tilde{\mathbf{g}}$. If $\zeta$ is uncorrelated with $\tilde{\mathbf{g}}$ then $\operatorname{Cov}(X,Y)=\operatorname{Cov}(X,Z)$ and $\operatorname{Var}(Y)=\operatorname{Var}(Z)+\sigma_{\zeta}^{2}$. Using
$\operatorname{Var}(X)$ $=\|\alpha\|_{2}^{2},$
$\operatorname{Cov}(X,Z)$ $=\alpha^{\top}\mathrm{sym}(B)\,\alpha,$
$\operatorname{Var}(Z)$ $\leq\|B\|_{2}^{2}\,\|\alpha\|_{2}^{2}.$
and the Rayleigh-Ritz principle, gives
$\rho_{XY}\ \geq\ \frac{\lambda_{\min}\!\big(\mathrm{sym}(B)\big)}{\sqrt{\ \|B\|_{2}^{2}+\sigma_{\zeta}^{2}/(\mathbf{u}_{\vartheta}^{\top}\Sigma_{g}\mathbf{u}_{\vartheta})\ }},$ (17)
which is Eq. (6) in the main content after identifying $B$. Without the uncorrelatedness assumption, Cauchy-Schwarz implies the fallback bound
$\rho_{XY}\ \geq\ \frac{\lambda_{\min}(\mathrm{sym}(B))\,\|\alpha\|_{2}-\sigma_{\zeta}}{\|B\|_{2}\,\|\alpha\|_{2}+\sigma_{\zeta}}.$ (18)
B.3 Negative Pair Curvature
In the projection-temperature subspace, we write $\mathbf{g}(z)=\mathbf{g}_{\vartheta}(z)$ and $\mathbf{u}=\mathbf{u}_{\vartheta}$. For softmax-type losses, a generalized Gauss-Newton curvature admits the population decomposition
$\mathbf{H}_{\vartheta}$ $=\Phi_{\text{pos}}+\Phi_{\text{neg}},$ (19)
$\Phi_{\text{pos}}$ $=\mathbb{E}_{z}\!\big[\mathbf{g}(z)\mathbf{g}(z)^{\top}\big],$
$\Phi_{\text{neg}}$ $=\mathbb{E}_{z\neq z^{\prime}}\!\big[\mathbf{g}(z)\mathbf{g}(z^{\prime})^{\top}\big],$
which mirrors the random-negative mechanism of InfoNCE. Note that $\Phi_{\text{pos}}\succeq 0$ while $\Phi_{\text{neg}}$ need not be PSD, and the sum $\mathbf{H}_{\vartheta}$ is PSD.
For a mini-batch $\{z_{i}\}_{i=1}^{B}$ let $\mathbf{g}_{i}=\mathbf{g}_{\vartheta}(z_{i})$, then
$\widehat{\Phi}_{\text{pos}}=\tfrac{1}{B}\sum_{i=1}^{B}\mathbf{g}_{i}\mathbf{g}_{i}^{\top},\hskip 28.80008pt\widehat{\Phi}_{\text{neg}}=\tfrac{1}{B(B-1)}\sum_{i\neq j}\mathbf{g}_{i}\mathbf{g}_{j}^{\top}.$ (20)
As in Eq. (8), define
$\mathbf{H}_{\vartheta}^{(\alpha)}=(1-\alpha)\,\Phi_{\text{pos}}+\alpha\,\Phi_{\text{neg}},\hskip 28.80008pt\mathbf{M}=\mathbf{H}_{\vartheta}^{(\alpha)}+\lambda\,\mathbf{I},$ (21)
with $\alpha\in[0,1]$ and $\lambda>0$. Let $A^{\star}(z)=\mathbf{g}(z)^{\top}\mathbf{H}_{\vartheta}^{-1}\mathbf{u}$, $A_{\alpha}(z)=\mathbf{g}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}$, and $\Delta_{\alpha}=\mathbf{H}_{\vartheta}-\mathbf{H}_{\vartheta}^{(\alpha)}$. Using the resolvent identity,
$\mathbf{M}^{-1}-\mathbf{H}_{\vartheta}^{-1}$ $=\mathbf{H}_{\vartheta}^{-1}\,(\Delta_{\alpha}-\lambda\mathbf{I})\,\mathbf{M}^{-1},$ (22)
$\big\|\mathbf{M}^{-1}-\mathbf{H}_{\vartheta}^{-1}\big\|_{2}$ $\leq\|\mathbf{H}_{\vartheta}^{-1}\|_{2}\,\|\Delta_{\alpha}\|_{2}\,\|\mathbf{M}^{-1}\|_{2}.$
which leads to
$\mathbb{E}_{z}\big[(A_{\alpha}-A^{\star})^{2}\big]\leq C_{2}\,\|\Delta_{\alpha}\|_{F}^{2}\,\|\mathbf{H}_{\vartheta}^{-1}\mathbf{u}\|_{2}^{2},$
for a constant $C_{2}>0$ depending only on $\|\Phi_{\text{pos}}\|_{2}$, $\|\mathbf{H}_{\vartheta}^{-1}\|_{2}$, and $\|\mathbf{M}^{-1}\|_{2}$. Whenever cross-example coupling is present, any $\alpha>0$ that injects off-diagonal mass reduces $\|\Delta_{\alpha}\|_{F}$. Let $\Pi_{k}\in\mathbb{R}^{k\times d_{\vartheta}}$ be a JL transform. For any fixed $\mathbf{a},\mathbf{b}$, $|\mathbf{a}^{\top}\mathbf{b}-(\Pi_{k}\mathbf{a})^{\top}(\Pi_{k}\mathbf{b})|\leq\varepsilon\|\mathbf{a}\|\,\|\mathbf{b}\|$ holds with probability at least $1-\delta$, where $k=\Omega(\varepsilon^{-2}\log(1/\delta))$. Define $\mathbf{M}_{k}=\Pi_{k}\,\mathbf{M}\,\Pi_{k}^{\top}$ and $\bar{\mathbf{u}}_{k}=\mathbf{M}_{k}^{-1}\Pi_{k}\mathbf{u}$. The sketched score
$\widehat{A}_{\alpha}(z)=(\Pi_{k}\mathbf{g}(z))^{\top}\bar{\mathbf{u}}_{k}$
satisfies $\mathbb{E}_{z}[(\widehat{A}_{\alpha}-A_{\alpha})^{2}]\leq C_{1}\log(1/\delta)/k$ for a constant $C_{1}>0$ depending on $\|\mathbf{M}^{-1}\|_{2}$, $\|\mathbf{u}\|_{2}$, and $\mathbb{E}\|\mathbf{g}(z)\|_{2}^{2}$. Combining with the mixing bias via the triangle inequality yields Eq. (9) in the main content:
$\mathbb{E}_{z}\!\big[\big(\widehat{A}_{\alpha}(z)-A^{\star}(z)\big)^{2}\big]$ $\leq\underbrace{\frac{C_{1}\log(1/\delta)}{k}}_{\text{projection variance}}$ (23)
$\quad+\underbrace{C_{2}\,\|\Delta_{\alpha}\|_{F}^{2}\,\|\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}\|_{2}^{2}}_{\text{curvature bias}}.$
Appendix C Data Selection Implementation Details
This section provides all components needed to reproduce CHIPS and baseline methods.
C.1 Problem Setup
Let the training pool be $\mathcal{D}_{\text{pool}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ and the evaluation set be $\mathcal{D}_{\text{eval}}=\{(x_{m},y_{m})\}_{m=1}^{M}$. Given a data retention ratio $r\in(0,1]$, we select $n=\lfloor r\times|\mathcal{D}_{\text{pool}}|\rfloor$ samples from $\mathcal{D}_{\text{pool}}$ to perform CPT. CHIPS operates in CLIP’s end-point subspace $\vartheta=\{\mathbf{W}_{v},\mathbf{W}_{t},\tau\}$ where $\mathbf{W}_{v}$ and $\mathbf{W}_{t}$ are the projection heads for vision and text, and $\tau>0$ is the logit scale. For computational efficienciy, backbone encoders are used to produce features and are kept fixed for computing selection utility scores (we only need training dynamics of $\vartheta$ instead of the complete model).
Let $\mathbf{h}=\mathrm{CLIP}_{\text{img}}(x)$ and $\mathbf{t}=\mathrm{CLIP}_{\text{txt}}(y)$ be backbone features. Define L2-normalized end-point embeddings
$\hat{\mathbf{x}}=\frac{\mathbf{W}_{v}\mathbf{h}}{\|\mathbf{W}_{v}\mathbf{h}\|},\hskip 28.80008pt\hat{\mathbf{y}}=\frac{\mathbf{W}_{t}\mathbf{t}}{\|\mathbf{W}_{t}\mathbf{t}\|},$
and similarities $s_{ij}=\tau\,\hat{\mathbf{x}}_{i}^{\top}\hat{\mathbf{y}}_{j}$. For a batch of size $B$, write $\mathbf{S}=[s_{ij}]_{i,j=1}^{B}$ and define the bidirectional softmax probabilities
$p^{i2t}_{ij}=\frac{\exp(s_{ij})}{\sum_{j^{\prime}}\exp(s_{ij^{\prime}})},\hskip 28.80008ptp^{t2i}_{ij}=\frac{\exp(s_{ij})}{\sum_{i^{\prime}}\exp(s_{i^{\prime}j})}.$
The symmetric InfoNCE loss for sample $i$ is
$\ell_{i}(\vartheta)=\tfrac{1}{2}\big(\mathrm{CE}(\mathbf{S}_{i,:},\,i)+\mathrm{CE}(\mathbf{S}_{:,i},\,i)\big),$
where $\mathrm{CE}$ denotes the cross-entropy loss.
Let $\mathbf{g}_{\vartheta}(z)=\nabla_{\vartheta}\ell(z;\vartheta)\in\mathbb{R}^{d}$ denote the end-point gradient of a sample, and let
$\mathbf{u}_{\vartheta}\ \triangleq\ \mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\vartheta}(z)]$
be the evaluation mean gradient in the same subspace. In practice we maintain $\mathbf{u}_{\vartheta}$ by an exponential moving average over random evaluation minibatches.
Moreover, to reduce computational cost we use a JL map matrix $\Pi_{k}\in\mathbb{R}^{k\times d}$ and work with sketched vectors $\mathbf{g}_{k}=\Pi_{k}\mathbf{g}_{\vartheta}$ and $\mathbf{u}_{k}=\Pi_{k}\mathbf{u}_{\vartheta}$. The inner product distortion satisfies $|\mathbf{a}^{\top}\mathbf{b}-(\Pi_{k}\mathbf{a})^{\top}(\Pi_{k}\mathbf{b})|\leq\varepsilon\|\mathbf{a}\|\,\|\mathbf{b}\|$ with probability at least $1-\delta$ when $k=\Omega(\varepsilon^{-2}\log(1/\delta))$, which is also used in App. B.3.
C.2 CHIPS
CHIPS ranks each $z\in\mathcal{D}_{\text{pool}}$ by the multiplicative utility
$\mathcal{I}_{\text{CHIPS}}(z)\ =\ \widehat{A}_{\alpha}(z)\cdot w_{\mathrm{L}}(z)\cdot w_{\mathrm{R}}(z),$
then selects the top $n$ samples for CPT. The three factors are implemented as follows and the complete process of CHIPS is provided in Alg. 1.
Curvature-aware proxy alignment.
We estimate the curvature in $\vartheta$ by mixing self and cross moments from symmetric InfoNCE, consistent with Sec. 2.3. Maintain EMA estimators
$\Phi_{\mathrm{pos}}=\mathbb{E}[\mathbf{g}_{\vartheta}(z)\mathbf{g}_{\vartheta}(z)^{\top}],\hskip 28.80008pt\Phi_{\mathrm{neg}}=\mathbb{E}_{z\neq z^{\prime}}[\mathbf{g}_{\vartheta}(z)\mathbf{g}_{\vartheta}(z^{\prime})^{\top}],$
then build
$\mathbf{M}\ =\ \mathbf{H}_{\vartheta}^{(\alpha)}+\lambda\mathbf{I},\hskip 28.80008pt\mathbf{H}_{\vartheta}^{(\alpha)}\ =\ (1-\alpha)\,\Phi_{\mathrm{pos}}+\alpha\,\Phi_{\mathrm{neg}},$
with $\alpha\in[0,1]$ and ridge $\lambda>0$. Without sketching, the alignment score is
$A_{\alpha}(z)\ =\ \mathbf{g}_{\vartheta}(z)^{\top}\,\mathbf{M}^{-1}\,\mathbf{u}_{\vartheta}.$
With a JL sketch, form $\mathbf{M}_{k}=\Pi_{k}\mathbf{M}\Pi_{k}^{\top}$ once per update and solve $\bar{\mathbf{u}}_{k}=\mathbf{M}_{k}^{-1}\mathbf{u}_{k}$. The sketched score is
$\widehat{A}_{\alpha}(z)\ =\ (\Pi_{k}\mathbf{g}_{\vartheta}(z))^{\top}\bar{\mathbf{u}}_{k}.$
This matches Eq. (4) and App. B.3.
Learnability.
Compute correctness and hardest negative margin using the current batch statistics
$p_{\mathrm{corr}}(z)$ $=\tfrac{1}{2}\big(p^{i2t}_{ii}+p^{t2i}_{ii}\big),$ (24)
$m(z)$ $=s_{ii}-\max\!\big\{\max_{j\neq i}s_{ij},\,\max_{i^{\prime}\neq i}s_{i^{\prime}i}\big\}.$
Define the learnability weight as in Eq. (11)
$w_{\mathrm{L}}(z)=\big(1-p_{\mathrm{corr}}(z)\big)\big(1+\sigma(-m(z))\big),$
which upweights near-boundary samples and downweights saturated ones.
Target-domain relevance.
Let the evaluation centroids be $\boldsymbol{\mu}_{x}=\mathbb{E}[\hat{\mathbf{x}}$] and $\boldsymbol{\mu}_{y}=\mathbb{E}[\hat{\mathbf{y}}$] over $\mathcal{D}_{\text{eval}}$. Following Eq. (12),
$w_{\mathrm{R}}(z)=\sigma\Big((1-\beta)\cos(\hat{\mathbf{x}},\boldsymbol{\mu}_{x})+\beta\cos(\hat{\mathbf{y}},\boldsymbol{\mu}_{y})\Big),$
which softly reweights toward the evaluation domain while keeping the selection drift bounded.
Algorithm 1 CHIPS Algorithm
1:pool $\mathcal{D}_{\text{pool}}$, evaluation set $\mathcal{D}_{\text{eval}}$, retention $n$, mix $\alpha$, ridge $\lambda$, JL dim $k$, relevance balance $\beta$
2:Estimate $\mathbf{u}_{\vartheta}\leftarrow\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\vartheta}(z)]$ using minibatches; cache $\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}$
3:Initialize EMA moments $\Phi_{\mathrm{pos}}\leftarrow\mathbf{0}$, $\Phi_{\mathrm{neg}}\leftarrow\mathbf{0}$
4:for each minibatch $\{z_{i}\}_{i=1}^{B}$ sampled from $\mathcal{D}_{\text{pool}}$ do
5:   compute $\mathbf{g}_{i}=\mathbf{g}_{\vartheta}(z_{i})$ for all $i$
6:   update $\Phi_{\mathrm{pos}}$ and $\Phi_{\mathrm{neg}}$ using the batch $U$-statistics and EMA
7:end for
[…]
9:if $k>0$ then
10:   sample JL $\Pi_{k}$, set $\mathbf{M}_{k}=\Pi_{k}\mathbf{M}\Pi_{k}^{\top}$, solve $\bar{\mathbf{u}}_{k}=\mathbf{M}_{k}^{-1}(\Pi_{k}\mathbf{u}_{\vartheta})$
11:end if
12:for each $z=(x,y)\in\mathcal{D}_{\text{pool}}$ do
13:   compute $\mathbf{g}=\mathbf{g}_{\vartheta}(z)$ and $\widehat{A}_{\alpha}(z)=\begin{cases}\mathbf{g}^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta},&k=0\\ (\Pi_{k}\mathbf{g})^{\top}\bar{\mathbf{u}}_{k},&k>0\end{cases}$
14:   compute $p_{\mathrm{corr}}(z)$ and $m(z)$, then $w_{\mathrm{L}}(z)=\big(1-p_{\mathrm{corr}}(z)\big)\big(1+\sigma(-m(z))\big)$
15:   compute $w_{\mathrm{R}}(z)=\sigma\big((1-\beta)\cos(\hat{\mathbf{x}},\boldsymbol{\mu}_{x})+\beta\cos(\hat{\mathbf{y}},\boldsymbol{\mu}_{y})\big)$
16:   set $\mathcal{I}_{\text{CHIPS}}(z)=\widehat{A}_{\alpha}(z)\,w_{\mathrm{L}}(z)\,w_{\mathrm{R}}(z)$
17:end for
18:return the top $n$ samples by $\mathcal{I}_{\text{CHIPS}}$
More details.
During the computation process,
  • $\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$ is reused across all samples, so the per-sample cost is dominated by $\mathbf{g}_{\vartheta}(z)$ and a dot product in $d$ or $k$ dimensions.
[…]
  • We rescale $\mathcal{I}_{\text{CHIPS}}$ within each shard of $\mathcal{D}_{\text{pool}}$ if memory constraints require sharded processing.
  • We recompute $\mathbf{u}_{\vartheta}$ periodically to track drift during CPT.
C.3 Heuristics-based Baselines
All heuristics select the top $n$ samples under their scores.
  • Random samples uniformly at random without replacement.
  • CLIPScore ranks by CLIPScore [21] computed with the frozen base CLIP model.
  • Concept-Balance performs probabilistic downsampling to flatten concept frequency. Following [36], we downsample overrepresented concepts such as Plots and Charts, Tables, and Scientific Formulae and Equations at rate $0.25$, then sample the remainder uniformly.
  • Concept-Filter keeps only samples whose metadata contain any of the eight whitelist concepts in [36], (Clinical Imaging, Microscopy, Immuno Assays, Illustrative Diagrams, Chemical Structures, Maps, Tools and Materials, and Hand Drawn and Screen Based Visuals) then samples uniformly from the filtered pool.
C.4 Influence-based Baselines in the End-point Subspace
For fairness and to avoid conflicts with the main text, all influence baselines use the same end-point gradients $\mathbf{g}_{\vartheta}$ and the same optional JL sketch $\Pi_{k}$ as CHIPS. Let $\tilde{\mathbf{g}}_{i}=\Pi_{k}\mathbf{g}_{\vartheta}(z_{i})$ if $k>0$ (use $\tilde{\mathbf{g}}_{i}=\mathbf{g}_{\vartheta}(z_{i})$ when $k=0$). Let the evaluation mean gradient be $\bar{\mathbf{g}}_{\text{eval}}=\frac{1}{M}\sum_{m=1}^{M}\tilde{\mathbf{g}}^{\,\text{eval}}_{m}$.
Dot [67].
The first-order directional alignment is
$\mathcal{I}_{\text{Dot}}(i)=\tilde{\mathbf{g}}_{i}^{\top}\bar{\mathbf{g}}_{\text{eval}}.$
TracIn [43].
Given checkpoints $\{\theta_{t}\}_{t=0}^{T-1}$ and learning rates $\{\eta_{t}\}$, we accumulate
$\mathcal{I}_{\text{TracIn}}(i)=\sum_{t=0}^{T-1}\eta_{t}\,\big(\tilde{\mathbf{g}}_{i}^{\,(t)}\big)^{\top}\bar{\mathbf{g}}_{\text{eval}},$
where $\tilde{\mathbf{g}}_{i}^{\,(t)}=\Pi_{k}\nabla_{\vartheta}\ell_{i}(\theta_{t})$. We keep $\bar{\mathbf{g}}_{\text{eval}}$ fixed for computational stability.
TRAK [40].
We form a regularized second-moment matrix in the same feature space
$\boldsymbol{\Phi}=\frac{1}{N}\sum_{i=1}^{N}\tilde{\mathbf{g}}_{i}\tilde{\mathbf{g}}_{i}^{\top}+\lambda_{\text{TRAK}}\mathbf{I},$
and compute
$\mathcal{I}_{\text{TRAK}}(i)=\tilde{\mathbf{g}}_{i}^{\top}\boldsymbol{\Phi}^{-1}\bar{\mathbf{g}}_{\text{eval}}.$
CHIPS for comparison.
CHIPS differs from Dot and TRAK by preconditioning with $\mathbf{M}$ that mixes $\Phi_{\text{pos}}$ and $\Phi_{\text{neg}}$ as in Sec. C.2, and by multiplying learnability and relevance weights. In the sketched space,
$\mathcal{I}_{\text{CHIPS}}(i)=\big(\tilde{\mathbf{g}}_{i}^{\top}\mathbf{M}_{k}^{-1}\mathbf{u}_{k}\big)\,w_{\mathrm{L}}(i)\,w_{\mathrm{R}}(i),$
where $\mathbf{M}_{k}=\Pi_{k}\big((1-\alpha)\Phi_{\mathrm{pos}}+\alpha\Phi_{\mathrm{neg}}+\lambda\mathbf{I}\big)\Pi_{k}^{\top}$.
Appendix D Score Distribution
The score distributions of CLIPScore, Dot, TracIn, TRAK, and CHIPS are provided in Fig. 5, Fig. 6, Fig. 7, Fig. 8, and Fig. 9.
Figure 5: Distribution of CLIPScore on BIOMEDICA.
Figure 6: Distribution of Dot on BIOMEDICA.
Figure 7: Distribution of TracIn on BIOMEDICA.
Figure 8: Distribution of TRAK on BIOMEDICA.
Figure 9: Distribution of CHIPS on BIOMEDICA.
Appendix E Training
E.1 Software and Distributed Setup
Framework and precision.
[…]
Communication.
Gradients are synchronized by NCCL with bucket size tuned to saturate NVLink. We enable gradient accumulation when the per-GPU micro-batch would otherwise exceed memory budget. Let $G$ be the number of GPUs, $B_{\text{micro}}$ the per-GPU micro-batch, and $A$ the accumulation steps, then the global batch is
$B_{\text{global}}=G\times B_{\text{micro}}\times A.$
We choose $G=8$, $B_{\text{micro}}=4096$ and $A=1$ such that $B_{\text{global}}=32{,}768$ as in the main content.
Determinism.
We fix seeds for Python, NumPy, and CUDA RNGs and set cuDNN to deterministic kernels where available. Dataset sharding and sampler seeds are logged per epoch to make re-runs bitwise reproducible up to nondeterminism in fused kernels.
E.2 Data Preprocessing and Batching
Image pipeline.
Training images are decoded to RGB and resized to the default input resolution of each MetaCLIP variant. We apply random resized crop that preserves aspect ratio, followed by horizontal flip with probability $0.5$. Pixel intensities are normalized by CLIP mean and variance. Center crop is used at evaluation time.
Text pipeline.
Texts are normalized by Unicode and then tokenized by the CLIP tokenizer. Sequences are truncated to the CLIP maximum length (77 tokens) with special tokens preserved.
Sharding and streaming.
The training pools are sharded into balanced files on disk to avoid hot spots. Each worker maintains a streaming iterator with prefetching in pinned memory. We keep sharding consistent across methods so that all runs process the same examples at the same step counts.
E.3 Optimization Schedule
Optimizer and parameter groups.
We use AdamW with $(\beta_{1},\beta_{2},\epsilon)=(0.9,0.98,10^{-6})$ as stated in the main text. Weight decay is applied to all trainable parameters except LayerNorm and bias parameters.
Learning-rate scheduler.
We use a cosine decay scheduler over the total number of optimizer steps. Unless specified otherwise, the peak learning rate equals the initial value $10^{-6}$. We optionally use a short linear warmup of $T_{\text{warm}}$ steps when training with heavier augmentation or larger $B_{\text{global}}$. The final learning rate floor is set to 0 unless explicit floor is reported.
E.4 Counting Steps and Budget Fairness
Let $n=\lfloor r\times|\mathcal{D}_{\text{train}}|\rfloor$ be the number of retained samples, $E$ the number of epochs, and $B_{\text{global}}$ the global batch size. The number of optimizer steps per run is
$T_{\text{steps}}\;=\;\left\lceil\frac{n}{B_{\text{global}}}\right\rceil\times E.$
All methods, including baselines and CHIPS, are trained for the same $T_{\text{steps}}$. The wall-clock overhead of selection is measured separately and reported as a percentage of the total training time.
Appendix F Evaluation
To ensure a comprehensive and reproducible evaluation, we include 48 datasets covering both general-domain and medical-domain visual understanding tasks for this paper. Detailed descriptions are provided below.
F.1 General-Domain Datasets
For General-Domain Datasets, they span a variety of recognition and retrieval tasks, including classification, fine-grained categorization, visual reasoning, robustness evaluation, and multimodal understanding. To be specific, we give an overall description for each as follows:
Object & Scene Classification:
  • ImageNet-1K [8]: A large-scale object classification benchmark with  1.28 million training images and 1,000 object categories, widely used as a baseline in deep vision.
  • ImageNetV2 [45]: A re-collected test set for ImageNet, designed to assess generalisation under dataset shift for the same 1,000 categories.
  • SUN397 [55]: A scene recognition dataset containing 397 scene categories of diverse indoor/outdoor types, enabling evaluation on scene-level classification.
  • Caltech-101 [13]: A mid-scale object classification benchmark with 101 categories and around 9,000 images, used historically for transfer-learning studies.
  • VOC2007 [12]: The PASCAL Visual Object Classes 2007 dataset, including object classification (and detection) tasks across 20 categories in real-world scenes.
Fine-Grained Recognition:
  • Cars [29]: A fine-grained vehicle classification dataset containing many car make/model/year classes, evaluating subtle visual differences among similar objects.
  • Aircraft [37]: A fine-grained aircraft dataset distinguishing among many aircraft models and variants, used for high-granularity recognition research.
  • Food101 [3]: A food image dataset covering 101 food types, with  101,000 images total, for fine-grained categorisation of dishes.
  • Oxford Flowers (Flowers102) [39]: A flower species classification dataset with 102 categories and  8,000 images, common in fine-grained vision tasks.
  • Pets [41]: A fine-grained pet breed classification dataset (cats & dogs) with multiple breeds and varied poses/backgrounds.
Texture, Shape, Small-Scale Objects:
  • CIFAR-10/100 [30]: Standard small-scale image classification datasets with 10 (CIFAR-10) and 100 (CIFAR-100) classes, each containing 60,000 colour images of 32×32 size.
  • MNIST [31]: Classic handwritten digit classification dataset with  70,000 grayscale images of digits 0-9, often used for benchmarking basic vision models.
  • STL-10 [7]: A dataset derived from ImageNet with 10 classes and high-resolution images, used for unsupervised / transfer learning in small-scale settings.
  • smallNORB (Azimuth/Elevation) [32]: A synthetic 3D object dataset capturing objects under different viewpoints (azimuth/elevation), for studying pose-invariance and representation robustness.
  • SVHN [38]: Street View House Numbers dataset with real-world digital number images, used for digit recognition in situ.
Robustness & Out-of-Distribution:
  • ImageNet-A [20]: A subset of ImageNet containing “adversarially filtered” images challenging standard models, designed to test worst-case object recognition robustness.
  • ImageNet-O [20]: An out-of-distribution test set for ImageNet models, containing images from unknown classes not present in training, to measure OOD detection/generalisation.
Domain-Specific Recognition:
  • KITTI [14]: A dataset collected for autonomous driving, including object, scene and motion tasks—here used in the classification context of road-scene recognition.
  • EuroSAT [19]: A remote sensing image classification dataset with 45 scene types, aimed at land-use and satellite-image recognition.
  • RESISC45 [5]: Another remote sensing scene classification dataset covering 45 classes of aerial images, evaluating models on high-altitude imagery.
Visual Reasoning & Synthetic Tasks:
  • CLEVR (Closest-Object-Distance / Count-All) [26]: A synthetic benchmark for visual reasoning, where models answer compositional questions about objects (distance/count) in rendered scenes.
  • DTD [6]: The Describable Texture Dataset, focusing on texture/material recognition with diverse patterns, supporting generalisation in less object-centric tasks.
Image-Text Retrieval / Multimodal:
  • Flickr8k [22]: An image-caption dataset with 8,000 images, used for evaluating image-to-text and text-to-image retrieval and captioning systems.
  • Flickr30k [63]: A larger image-caption dataset with 30,000 images, used widely in cross-modal retrieval research.
  • MSCOCO [33]: A large-scale multimodal dataset with 120K+ images and captions, supporting image detection, segmentation and retrieval tasks.
  • Rendered-SST2 [44]: A dataset created by rendering the sentences from the Stanford Sentiment Treebank v2 into images (positive/negative labels), used to assess optical-character-recognition via image encoders. (Train: 6,920 images; Val: 872; Test: 1,821).
Geographic / Landmark Classification:
  • Country211 [44]: A dataset designed for geolocation classification, filtered from the YFCC100M dataset—211 countries/territories, with 150 train / 50 val / 100 test images per country.
F.2 Medical-Domain Datasets
For Medical-Domain Datasets, we include a broad range of imaging modalities and clinical specialties, covering ophthalmology, radiology, dermatology, hematology, pathology, neuropathology, and non-clinical biology. These datasets collectively assess model performance in clinically relevant scenarios and diagnostic contexts. From a task perspective, they can be broadly grouped into three major categories: disease classification, organ and tissue recognition, and pathological image analysis, with an additional category for non-clinical biological imaging. Together, they form a comprehensive benchmark for evaluating generalization and robustness of visual models in biomedical applications. In detail, the brief introduction of included datasets are listed:
Disease Classification:
  • Diabetic Retinopathy [11]: A fundus-image dataset for grading diabetic retinopathy severity in ophthalmology, used for multi-class disease classification.
  • RetinaMNIST / OCTMNIST [61]: Retinal fundus and optical coherence tomography (OCT) image datasets for ophthalmic disease classification across multiple categories; part of the MedMNIST benchmark (about 708k 2D images in total for the MedMNIST collection).
  • ChestMNIST [61]: Based on the NIH-ChestXray14 dataset, ChestMNIST contains 112,120 chest X-ray images, formulated as a multi-label classification task for detecting 14 different diseases.
  • ChestX-ray14 [52]: A large collection of over 100,000 frontal-view chest X-ray images from more than 30,000 patients. The labels for eight common thoracic diseases were automatically generated by text-mining the associated radiological reports.
  • DermaMNIST [61]: This is a multi-class classification dataset of 10,015 dermatoscopic images for identifying 7 different types of common pigmented skin lesions.
  • BloodMNIST [61]: A hematology microscope-image dataset containing 17,092 images across 8 classes, used for blood-cell type classification.
Organ and Tissue Recognition:
  • OrganAMNIST / OrganCMNIST / OrganSMNIST [61]: This dataset is for multi-class classification of 11 body organs, containing 58,850 images derived from axial-view slices of abdominal CT scans and resized to 28x28 pixels.
  • TissueMNIST [61]: Sourced from the Broad Bioimage Benchmark Collection, TissueMNIST is a large-scale dataset of 236,386 human kidney cortex cells, organized into 8 categories for a multi-class classification task.
Pathological Image Analysis:
  • PCAM [51]: The dataset is a large-scale collection of 96x96 pixel histopathology image patches extracted from the Camelyon16 challenge, designed for the task of identifying metastatic cancer in lymph node sections.
  • LC25000 [2]: The LC25000 dataset contains 25,000 color histopathological images across five classes, featuring both cancerous and benign tissues from the lung and colon.
  • PathMNIST [61]: PathMNIST is a multi-class classification dataset derived from colorectal cancer histology slides. It is comprised of 107,180 image patches, categorized into 9 distinct tissue types.
  • Amyloid CAA/Diffuse [53]: It includes 100495 annotations on 20099 candidate amyloid beta neuropathologies, which is a neuropathology image dataset for subtypes of amyloid pathology in brain tissue, used for subtype classification tasks.
Non-Clinical Biological Imaging:
  • Pollen [1]: The Pollen13K dataset is a large-scale collection of over 13,000 microscopic pollen grain images from aerobiological samples, used for biological particle classification. It was chosen to test model generalization on complex, non-clinical imagery.
Appendix G FLOPs Computation
This appendix consolidates the FLOPs counting procedure for BIOMEDICA [36] used to report scoring cost.
All totals below measure a single complete scoring pass over $\mathcal{D}_{\text{train}}$ and use the batch primitives in Tab. 5: $C_{\text{fwd}}(B)$, $C_{\text{bwd}}(B)$, $C_{\text{fb}}(B)=C_{\text{fwd}}(B)+C_{\text{bwd}}(B)$, and $C_{\text{jvp}}(B)$. The random projection (CountSketch) cost per application is $C_{\text{rp}}\approx 2P$ (see Sec. G). We denote $n_{\text{train}}=\lceil N_{\text{train}}/B_{\text{train}}\rceil$ and $n_{\text{eval}}=\lceil N_{\text{eval}}/B_{\text{eval}}\rceil$. TracIn uses $E$ epochs for accumulation. TRAK and CHIPS use $I$ conjugate-gradient (CG) iterations.
TracIn.
TracIn combines a single evaluation-direction construction with per-batch JVPs and trajectory accumulation:
$C_{\text{TracIn}}$ $=C_{\text{eval-dir}}+C_{\text{JVP-train}}+C_{\text{accum}},$ (25)
$C_{\text{eval-dir}}$ $=n_{\text{eval}}\,\big(C_{\text{fb}}(B_{\text{eval}})+C_{\text{rp}}\big),$ (26)
$C_{\text{JVP-train}}$ $=n_{\text{train}}\,C_{\text{jvp}}(B_{\text{train}}),$ (27)
$C_{\text{accum}}$ $=E\,n_{\text{train}}\,C_{\text{fb}}(B_{\text{train}}).$ (28)
TRAK.
TRAK builds a second-order score with CG in the same end-point geometry. We write the shared backbone block once and reuse it:
$\mathrm{Base}$ $=n_{\text{eval}}\,\big(C_{\text{fb}}(B_{\text{eval}})+C_{\text{rp}}\big)\;+\;C_{\text{proto-eval}}$
$\quad+\;n_{\text{train}}\,\big(C_{\text{fb}}(B_{\text{train}})+C_{\text{rp}}\big)$
$\quad+\;I\,n_{\text{train}}\,\big(C_{\text{jvp}}(B_{\text{train}})+C_{\text{fb}}(B_{\text{train}})+C_{\text{rp}}\big)$
$\quad+\;n_{\text{train}}\,\big(C_{\text{jvp}}(B_{\text{train}})+C_{\text{fwd}}(B_{\text{train}})\big).$ (29)
The TRAK total is simply
$C_{\text{TRAK}}=\mathrm{Base}.$ (30)
Quantity Meaning General formula $\boldsymbol{B_{\text{train}}}$ $\boldsymbol{B_{\text{eval}}}$
$C_{\text{lin}}$ two linear projections to $d$ $2\,B\,(d_{v}{+}d_{t})\,d$ $294967296\times 10^{10} 4.4$$456448\times 10^{9}$
$C_{\text{norm}}$ two L2 normalizations $\approx 6\,B\,d$ $1.00663296\times 10^{8}$ $1.04448\times 10^{7}$
$C_{\text{mm}}$ logits matmul for one direction $2\,B^{2}\,d$ $099511627776\times 10^{12} 1.1$$183744\times 10^{10}$
$C_{\text{fwd}}$ forward, both directions $C_{\text{lin}}+C_{\text{norm}}+2C_{\text{mm}}$ $2.242073591808\times 10^{12}$ $2.81417728\times 10^{10}$
$C_{\text{bwd}}$ backward to end-point heads and $\tau$ $\approx 2\,(C_{\text{lin}}+2C_{\text{mm}})$ $4.483945857024\times 10^{12}$ $5.6262656\times 10^{10}$
$C_{\text{fb}}$ forward + backward $C_{\text{fwd}}+C_{\text{bwd}}$ $6.726019448832\times 10^{12}$ $8.44044288\times 10^{10}$
$C_{\text{jvp}}$ JVP cost upper bound $\approx 2\,C_{\text{fwd}}$ $4.484147183616\times 10^{12}$ $5.62835456\times 10^{10}$
$C_{\text{proto-eval}}$ eval prototypes only $C_{\text{lin}}(B_{\text{eval}})+C_{\text{norm}}(B_{\text{eval}})$ $4.4668928\times 10^{9}$
Table 5: Batch-level FLOPs primitives and instantiated costs. Symmetric CLIP loss computes logits in both directions. $C_{\text{jvp}}$ uses the empirical bound $C_{\text{jvp}}\approx 2\,C_{\text{fwd}}$ for this architecture and batching.
CHIPS.
CHIPS shares $\mathrm{Base}$ with TRAK and adds three lightweight terms for (i) negative-pair curvature in the sketched space, (ii) hardest-negative margin bookkeeping, and (iii) two prototype similarities per batch for relevance. To avoid overlong lines we factor these contributions:
$\Delta_{\text{neg}}$ $\approx c_{\text{neg}}\,k$ (31)
$\Delta_{\text{margin}}$ $\approx 2\,B_{\text{train}}^{2}$ (32)
$\Delta_{\text{rel}}$ $\approx 4\,B_{\text{train}}\,d$ (33)
which are vector-level ops per CG iteration, hardest-negative search in a $B_{\text{train}}\times B_{\text{train}}$ block, and two prototype similarities per batch.
The CHIPS total is then
$C_{\text{CHIPS}}$ $=\mathrm{Base}$
$\quad+\;I\,\Delta_{\text{neg}}$
$\quad+\;n_{\text{train}}\,\Delta_{\text{margin}}$
$\quad+\;n_{\text{train}}\,\Delta_{\text{rel}}.$ (34)
$C_{\text{rp}}$ is incurred when forming the sketched evaluation direction and wherever sketched vectors are refreshed. $\Delta_{\text{neg}}$ accounts for a handful of axpy-like operations per CG iteration in the $k$-dimensional sketched space. $\Delta_{\text{margin}}$ is a max-reduction over off-diagonal logits already materialized for symmetric InfoNCE. $\Delta_{\text{rel}}$ computes two dot products per sample with cached prototypes $(\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y})$.
G.1 Numerical Totals
We instantiate Eqs. (25)–(34) with the fixed values above. We use $E=10$ epochs for TracIn and $I=5$ CG iterations for TRAK and CHIPS.
$\textbf{TracIn}\ (E{=}10):\quad$ $5.258869\times 10^{16}.$
$\textbf{TRAK}\ (I{=}5):\quad$ $5.094585\times 10^{16}.$
$\textbf{CHIPS}\ (I{=}5):\quad$ $5.094747\times 10^{16}.$
G.2 Assumptions and Further notes
  • Symmetric CLIP computes logits in both directions; $C_{\text{fwd}}(B)$ already includes the two $B\times B$ matmuls in Eq. (5).
  • The empirical bound $C_{\text{jvp}}(B)\approx 2\,C_{\text{fwd}}(B)$ holds for our implementation and batch shapes.
  • The evaluation mean-gradient uses a single split with $B_{\text{eval}}=3400$. $C_{\text{proto-eval}}$ counts only projections and L2 norms.
  • CountSketch is the default random projection; for other sketches replace $C_{\text{rp}}$ with the appropriate cost model (sparse RP $2sP$, SRHT $2m\log_{2}m$, dense Gaussian $2kP$).
  • FLOPs are operation counts independent of arithmetic precision and exclude file I/O and host preprocessing.
  • Scaling summary. With fixed $B$ and $d$, the dominant terms scale as TracIn $\Theta(E\,n_{\text{train}}\,C_{\text{fb}})$, TRAK $\Theta(I\,n_{\text{train}}\,C_{\text{fb}})$, CHIPS $\Theta(I\,n_{\text{train}}\,C_{\text{fb}})+O(Ik+n_{\text{train}}B^{2}+n_{\text{train}}Bd)$. At the reported $k$ and $B$, the $O(Ik)$ and $O(n_{\text{train}}Bd)$ extras are negligible and $\Delta_{\text{margin}}$ is amortized by already-computed logits.
Appendix H Full Results
H.1 Main Experiment
The full results of the main experiment are detailed in Tab. 6 and Tab. 7 for the medical domain, and in Tab. 8 through Tab. 11 for the general domain.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
PubMedCLIP 63.25 51.46 8.16 25.27 94.12 26.88 8.45 19.01
BioMedCLIP 2.26 47.53 20.19 72.91 1.65 87.92 10.14 3.43
BMCLIP 69.95 73.96 38.00 17.28 1.80 74.72 9.44 8.77
Vanilla 2.58 55.77 20.15 10.56 52.61 63.31 17.95 57.67
Full Dataset 60.89 71.62 44.83 9.81 88.30 77.27 23.33 62.11
rr=10%
Random 67.29 68.43 38.91 10.18 19.85 68.07 20.61 8.44
Concept-Balance 61.18 66.89 39.71 10.87 11.07 82.57 21.81 15.62
Concept-Filter 54.91 68.60 40.00 11.06 6.85 82.21 23.41 9.93
CLIPScore 2.49 60.68 40.67 10.56 72.81 55.49 20.99 27.80
Dot 71.67 57.20 39.89 23.51 72.64 18.34 16.87 5.50
TracIn 65.40 56.75 40.03 26.34 90.73 12.84 16.69 3.69
TRAK 43.23 59.23 39.57 12.88 77.16 14.74 25.93 3.60
CHIPS (ours) 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
rr=20%
Random 65.86 57.13 40.55 11.13 10.30 62.52 21.92 15.43
Concept-Balance 55.50 69.32 40.13 10.37 3.49 82.83 24.00 15.01
Concept-Filter 54.92 69.78 41.04 15.21 3.25 68.44 25.46 6.93
CLIPScore 4.34 59.92 47.44 11.25 48.38 16.31 18.82 11.17
Dot 72.11 60.07 31.49 33.63 62.01 45.43 13.30 11.07
TracIn 72.77 57.24 36.61 19.17 75.28 16.98 19.70 3.00
TRAK 71.76 58.73 41.71 28.16 42.51 15.15 16.75 3.31
CHIPS (ours) 72.50 68.67 48.07 30.36 53.58 23.50 22.49 13.97
rr=30%
Random 70.95 70.79 41.39 10.94 29.82 64.34 18.42 14.17
Concept-Balance 55.50 62.83 41.01 10.43 5.07 66.69 24.61 10.15
Concept-Filter 54.92 52.14 41.15 21.06 5.73 76.31 25.96 3.97
CLIPScore 16.89 53.70 44.16 11.94 29.20 23.78 13.94 5.59
Dot 70.35 57.56 32.45 20.74 44.69 25.38 15.40 6.39
TracIn 72.13 58.48 29.81 22.19 50.72 31.45 14.24 7.28
TRAK 69.39 59.76 42.11 25.14 35.00 16.40 16.92 1.64
CHIPS (ours) 73.42 73.66 49.88 36.34 67.90 25.15 25.02 14.39
rr=50%
Random 63.57 72.99 39.65 10.56 12.41 69.59 18.82 17.01
Table 6: Full medical-domain results (Part 1/2) of the main experiment.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
PubMedCLIP 12.83 9.63 26.50 23.07 22.63 23.81 21.73 18.00 8.45
BioMedCLIP 9.78 1.20 25.00 20.28 22.55 21.92 8.48 16.00 4.66
BMCLIP 9.57 61.65 25.70 27.15 20.37 21.35 43.52 19.75 5.02
Vanilla 9.35 10.82 26.40 10.42 9.91 11.46 22.21 18.75 5.24
Full Dataset 8.64 11.72 21.50 17.51 14.89 14.20 40.91 17.50 12.40
rr=10%
Random 6.57 11.27 23.00 13.80 10.35 10.49 25.75 17.75 4.68
Concept-Balance 7.63 11.67 21.60 15.07 11.14 11.09 27.49 16.75 5.03
Concept-Filter 7.44 11.72 22.30 14.79 10.43 11.46 31.98 17.00 5.14
CLIPScore 6.86 11.47 27.20 8.51 8.99 11.60 25.79 16.75 5.25
Dot 6.17 12.17 20.60 9.48 10.36 11.19 25.42 33.75 4.16
TracIn 6.53 12.02 18.00 13.98 12.85 15.28 22.98 40.50 4.96
TRAK 25.84 12.17 20.90 10.74 10.43 11.22 28.68 26.75 4.46
CHIPS (ours) 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
rr=20%
Random 17.84 11.62 24.00 14.75 10.13 10.93 29.86 17.50 6.60
Concept-Balance 9.46 11.97 25.30 17.65 13.19 12.77 33.75 17.50 5.42
Concept-Filter 8.41 11.97 22.40 17.28 13.01 12.85 34.54 16.75 6.42
CLIPScore 4.94 11.27 27.40 10.10 11.25 11.69 22.34 20.25 4.45
Dot 7.15 11.77 22.00 12.35 12.50 12.87 25.14 23.50 5.38
TracIn 11.96 11.82 17.40 13.39 13.94 19.12 28.91 37.25 5.07
TRAK 6.94 11.02 23.50 12.42 13.49 14.43 25.10 34.50 5.41
CHIPS (ours) 8.96 11.87 21.40 13.50 13.80 14.87 23.20 38.50 5.92
rr=30%
Random 7.93 12.07 22.10 15.32 10.76 11.33 30.60 17.25 6.42
Concept-Balance 8.40 12.07 23.60 16.53 12.38 11.92 33.11 16.75 6.48
Concept-Filter 7.52 12.02 25.20 17.38 13.72 13.57 36.16 17.00 6.81
CLIPScore 4.19 11.32 26.90 9.64 11.16 10.69 28.69 22.00 4.49
Dot 10.49 11.62 21.40 13.21 12.61 12.56 30.03 23.50 7.69
TracIn 17.75 11.97 18.00 14.57 14.57 16.74 33.68 28.75 5.69
TRAK 6.49 11.42 19.20 13.80 14.61 15.91 22.98 28.25 7.75
CHIPS (ours) 10.65 11.37 21.30 13.13 14.28 15.33 24.43 30.25 7.69
rr=50%
Random 9.86 11.82 22.60 16.86 12.66 12.90 35.31 17.25 10.06
Table 7: Full medical-domain results (Part 2/2) of the main experiment. Abbreviations: Derma = DermaMNIST, Oct = OctMNIST, OrganA = OrganAMNIST, OrganC = OrganCMNIST, OrganS = OrganSMNIST, Path = PathMNIST, Retina = RetinaMNIST, Tissue = TissueMNIST.
Model Cars Country211 FER Aircraft Food101 GTSRB
Imagenet
-A
Imagenet
-O
Imagenet
1k
Imagenet
v2
PubMedCLIP 8.51 3.09 20.30 4.32 15.43 15.37 4.97 14.55 14.95 13.00
BioMedCLIP 0.37 0.37 13.33 1.02 0.99 2.49 0.72 0.75 0.07 0.07
BMCLIP 92.13 27.18 37.57 31.38 90.40 56.22 54.93 33.85 71.97 64.80
Vanilla 74.29 22.41 42.57 28.50 87.24 43.80 46.97 39.55 70.78 62.59
Full Dataset 66.11 20.41 45.00 20.37 81.00 36.94 42.24 37.60 66.28 58.90
rr=10%
Random 71.43 21.76 43.73 27.06 84.99 41.94 45.60 39.20 69.18 61.31
Concept-Balance 71.28 21.81 44.90 27.48 84.82 42.76 45.73 39.35 69.12 61.05
Concept-Filter 70.92 21.61 43.86 26.91 84.59 42.36 44.91 39.30 68.75 60.85
CLIPScore 74.93 22.42 42.09 28.77 86.90 44.25 47.16 39.65 70.60 62.63
Dot 65.32 17.74 38.12 19.98 76.25 29.78 37.57 31.10 60.91 53.14
TracIn 64.82 17.30 35.94 19.71 74.90 27.40 36.41 29.95 59.22 51.16
TRAK 65.10 17.69 37.46 20.58 76.27 28.89 37.61 30.15 60.54 52.58
CHIPS (ours) 65.61 17.80 38.72 20.43 75.94 29.65 38.13 30.25 61.07 53.01
rr=20%
Random 70.20 21.60 45.29 25.71 84.47 41.35 45.23 39.30 68.67 60.53
Concept-Balance 69.12 21.29 45.89 25.77 83.75 39.30 44.53 38.40 68.26 60.30
Concept-Filter 69.42 21.20 44.61 25.89 83.36 38.61 44.03 38.40 67.87 59.94
CLIPScore 71.88 20.70 41.32 26.79 85.41 41.09 45.59 35.95 68.13 60.13
Dot 70.84 18.83 39.24 22.86 81.28 29.65 40.45 31.75 63.04 55.01
TracIn 69.54 18.55 39.29 22.23 80.37 29.30 39.48 30.55 62.51 54.46
TRAK 69.44 18.79 37.98 23.34 80.05 31.58 41.32 31.85 63.56 55.44
CHIPS (ours) 69.32 19.54 39.40 24.36 80.05 31.06 40.33 32.10 62.97 55.01
rr=30%
Random 69.36 21.36 46.22 25.23 84.10 39.74 44.61 38.40 68.21 60.20
Concept-Balance 69.03 21.50 46.20 25.92 83.55 38.97 44.56 38.60 68.33 60.48
Concept-Filter 68.88 20.82 43.81 25.08 82.21 36.93 42.68 37.75 67.13 59.44
CLIPScore 72.11 20.48 42.05 26.46 84.78 38.42 44.92 35.40 67.68 59.30
Dot 69.36 18.16 37.94 22.47 79.58 27.87 39.15 30.40 61.90 53.53
TracIn 68.96 18.20 39.84 22.89 79.51 29.26 39.65 30.15 61.87 53.63
TRAK 68.82 18.36 38.63 23.91 79.20 30.37 39.47 30.75 62.41 54.55
CHIPS (ours) 68.30 18.39 37.89 23.46 79.27 30.36 38.96 31.40 62.33 54.45
rr=50%
Random 68.47 21.03 46.15 23.28 82.96 38.74 43.67 37.90 67.61 59.90
Table 8: Full general-domain results (Part 1/4) of the main experiment.
Model MNIST
Rendered
SST2
STL10 Sun397
Sun397
Official
VOC Caltech101 CIFAR10 CIFAR100
PubMedCLIP 25.55 54.37 59.24 18.27 16.86 34.95 22.73 38.40 10.54
BioMedCLIP 9.58 50.08 9.70 0.36 0.24 3.55 0.51 10.31 1.19
BMCLIP 86.27 63.92 97.86 69.79 69.28 80.87 83.55 96.98 81.53
Vanilla 47.83 60.57 97.28 66.83 68.20 72.18 83.80 90.24 66.69
Full Dataset 27.61 55.24 97.29 65.17 65.64 66.12 83.73 91.43 66.79
rr=10%
Random 45.07 55.85 97.05 66.82 67.89 69.95 85.55 91.94 67.61
Concept-Balance 37.41 54.26 97.12 67.01 68.09 69.44 85.85 92.34 68.40
Concept-Filter 41.63 56.29 96.96 66.88 67.80 69.87 85.55 92.25 67.94
CLIPScore 47.82 60.57 97.28 66.82 68.45 72.70 84.40 90.83 67.90
Dot 38.78 55.63 95.54 58.91 61.10 68.78 81.79 85.03 58.04
TracIn 39.09 52.94 95.24 57.33 59.55 68.56 82.17 82.60 54.60
TRAK 44.99 56.84 95.80 58.33 60.34 66.35 83.22 82.17 54.39
CHIPS (ours) 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
rr=20%
Random 34.08 54.48 97.17 66.71 67.57 67.92 85.39 92.22 67.98
Concept-Balance 38.87 55.68 97.14 66.54 67.39 67.91 85.09 91.96 67.62
Concept-Filter 36.40 57.11 96.88 66.43 66.99 69.49 85.19 91.68 67.26
CLIPScore 47.78 58.54 97.04 62.08 64.06 71.39 83.63 88.03 61.89
Dot 38.50 57.77 95.36 56.67 59.19 66.22 81.61 84.61 57.69
TracIn 37.95 55.19 94.67 56.22 58.38 67.15 81.71 84.93 56.99
TRAK 43.87 51.62 95.14 55.84 58.63 68.10 82.04 85.28 58.14
CHIPS (ours) 42.91 50.63 95.42 55.94 58.51 65.89 81.38 85.72 59.00
rr=30%
Random 27.07 54.26 97.15 66.51 67.18 67.28 85.08 92.34 68.13
Concept-Balance 31.72 54.31 97.00 66.68 67.37 68.08 85.09 92.18 68.32
Concept-Filter 34.15 57.33 96.80 65.84 66.55 68.56 84.34 91.07 65.50
CLIPScore 46.31 57.06 97.05 62.11 63.81 69.97 83.32 88.35 62.38
Dot 37.90 54.64 95.31 55.82 57.89 66.75 80.82 83.64 55.78
TracIn 37.04 53.21 94.96 55.72 57.92 67.11 80.84 83.05 54.84
TRAK 32.68 50.08 95.30 55.68 57.78 64.52 80.81 84.56 57.45
CHIPS (ours) 34.25 49.92 95.36 55.69 57.91 64.36 80.79 84.61 57.64
rr=50%
Random 30.17 54.09 97.21 66.02 66.73 66.76 84.60 92.11 67.94
Table 9: Full general-domain results (Part 2/4) of the main experiment.
Model
CLEVR
Closest
CLEVR
Count
DMLAB DTD Eurosat Flowers KITTI
PubMedCLIP 22.28 16.23 18.85 10.59 19.70 14.54 31.36
BioMedCLIP 24.51 16.85 16.86 1.17 11.26 0.59 29.54
BMCLIP 15.79 33.98 13.80 55.64 63.02 76.11 26.30
Vanilla 22.47 29.11 16.11 56.22 55.96 73.57 24.47
Full Dataset 20.45 21.03 12.08 50.48 50.06 70.69 32.49
rr=10%
Random 22.57 29.39 12.28 53.56 53.31 72.68 31.08
Concept-Balance 22.59 31.12 12.02 52.93 52.39 72.94 28.83
Concept-Filter 22.41 31.32 11.99 52.66 48.26 72.68 29.54
CLIPScore 22.45 31.71 15.08 56.60 56.54 74.21 22.93
Dot 21.38 25.25 15.86 47.82 42.61 66.14 17.58
TracIn 21.39 22.96 14.83 45.80 40.17 66.08 17.02
TRAK 22.55 21.23 14.04 48.14 37.98 68.12 20.82
CHIPS (ours) 22.59 21.31 14.92 48.86 40.27 66.29 19.83
rr=20%
Random 22.16 26.30 11.98 52.66 51.19 73.05 30.66
Concept-Balance 21.53 29.22 11.80 52.07 52.26 72.48 29.68
Concept-Filter 21.51 30.64 11.92 51.91 50.33 71.74 31.50
CLIPScore 22.47 28.19 14.92 55.00 48.13 70.60 20.39
Dot 20.93 20.89 16.73 46.44 41.33 65.67 20.39
TracIn 18.71 21.03 15.45 43.88 40.69 64.55 16.60
TRAK 21.21 24.05 15.61 45.11 41.83 66.45 18.71
CHIPS (ours) 21.52 25.25 16.21 44.47 42.89 66.53 16.46
rr=30%
Random 21.86 26.93 11.85 52.34 52.33 72.56 32.07
Concept-Balance 21.23 27.33 11.83 51.44 52.69 72.56 30.38
Concept-Filter 21.63 27.87 11.84 51.65 48.91 71.59 31.50
CLIPScore 22.44 27.65 14.88 55.11 47.78 70.92 18.00
Dot 21.01 20.81 15.41 42.82 39.65 64.90 16.60
TracIn 18.77 20.52 15.44 43.14 37.85 65.05 17.02
TRAK 21.40 20.57 15.64 44.26 41.35 64.68 15.61
CHIPS (ours) 21.60 20.78 15.76 44.20 42.19 65.30 14.77
rr=50%
Random 21.29 24.69 11.74 51.38 50.54 71.85 32.77
Table 10: Full general-domain results (Part 3/4) of the main experiment.
Model Pets RESISC45
Smallnorb
Azimuth
Smallnorb
Elevation
SVHN
PubMedCLIP 23.36 14.22 5.51 10.97 7.43
BioMedCLIP 2.89 2.21 5.64 10.95 9.83
BMCLIP 91.39 63.38 6.29 10.54 50.43
Vanilla 90.54 66.08 5.42 10.92 24.34
Full Dataset 88.77 63.05 6.77 12.12 19.39
rr=10%
Random 90.52 64.08 5.93 10.88 18.41
Concept-Balance 90.11 63.76 5.82 10.47 19.17
Concept-Filter 89.83 62.35 5.82 10.53 19.56
CLIPScore 90.02 66.38 5.67 11.51 25.80
Dot 88.09 60.21 5.84 10.44 26.41
TracIn 88.23 60.14 5.47 11.37 24.96
TRAK 87.57 60.54 5.26 11.60 26.31
CHIPS (ours) 86.97 58.44 5.55 11.82 25.23
rr=20%
Random 90.16 63.87 6.11 10.97 18.19
Concept-Balance 89.45 63.49 6.07 11.05 18.78
Concept-Filter 89.32 61.79 5.90 11.12 20.29
CLIPScore 89.02 64.83 5.34 10.80 27.57
Dot 86.48 60.22 6.51 10.48 25.77
TracIn 88.25 59.75 5.37 10.96 25.78
TRAK 86.59 60.29 5.67 11.80 24.32
CHIPS (ours) 87.24 59.68 5.49 11.02 24.63
rr=30%
Random 89.92 63.79 6.27 10.74 17.19
Concept-Balance 89.29 64.56 6.41 10.78 18.74
Concept-Filter 89.23 62.60 5.97 11.74 20.28
CLIPScore 88.39 64.11 5.13 10.89 26.40
Dot 86.59 60.65 5.69 11.64 23.29
TracIn 87.76 59.51 5.54 10.67 23.28
TRAK 86.84 59.44 5.60 11.30 23.07
CHIPS (ours) 87.08 60.00 5.84 11.18 22.45
rr=50%
Random 89.64 63.71 6.19 11.04 17.09
Table 11: Full general-domain results (Part 4/4) of the main experiment.
H.2 Generalization Experiment
The full results of the main experiment are detailed in Tab. 12 and Tab. 13 for the medical domain, and in Tab. 14 through Tab. 17 for the general domain.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
B32-400M
Random 53.80 65.05 40.88 67.57 93.21 23.61 14.56 1.24
TracIn 57.38 55.25 37.12 50.47 96.33 12.32 21.66 2.16
CHIPS (ours) 59.94 55.22 36.27 69.83 98.35 13.97 16.22 1.56
B32-$C$
Random 73.20 62.22 21.31 32.24 1.54 87.94 35.52 3.53
TracIn 66.56 56.44 19.17 23.26 5.01 85.16 30.20 11.05
CHIPS (ours) 51.66 56.75 19.44 39.85 6.46 87.88 28.73 34.37
B16-400M
Random 67.29 68.43 38.91 10.18 19.85 68.07 20.61 8.44
TracIn 65.40 56.75 40.03 26.34 90.73 12.84 16.69 3.69
CHIPS (ours) 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
B16-$C$
Random 72.85 59.01 20.61 16.66 49.21 87.68 16.90 2.33
TracIn 74.49 52.07 20.33 19.04 91.57 37.16 12.36 2.00
CHIPS (ours) 76.91 54.99 32.75 22.88 34.65 82.59 14.88 2.90
L14-400M
Random 73.65 55.76 43.68 19.61 11.95 87.16 9.97 5.72
TracIn 65.62 60.57 40.61 16.03 18.51 51.40 19.47 5.10
CHIPS (ours) 65.34 61.27 39.65 19.48 44.86 73.60 11.14 5.84
L14-$C$
Random 47.71 54.50 24.16 68.13 11.29 86.15 18.06 41.76
TracIn 51.44 56.73 26.40 60.72 42.54 83.83 18.80 49.22
CHIPS (ours) 54.22 56.09 21.73 63.54 33.75 84.75 20.50 46.61
H14-$C$
Random 72.41 64.77 50.27 12.19 92.56 86.96 24.44 2.43
TracIn 73.09 65.87 50.75 12.70 76.15 83.24 14.24 2.92
CHIPS (ours) 78.64 65.30 55.12 11.50 82.87 86.85 20.94 8.00
Table 12: Full medical-domain results (Part 1/2) of the generalization experiment.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
B32-400M
Random 6.05 13.97 25.00 10.77 6.86 5.30 34.50 9.75 5.31
TracIn 3.80 17.11 24.80 9.93 10.32 9.19 16.56 36.00 12.49
CHIPS (ours) 2.64 12.52 24.20 13.31 11.33 10.59 27.87 19.25 11.63
B32-$C$
Random 4.42 9.63 25.00 16.22 8.98 6.24 31.91 14.25 9.18
TracIn 6.71 16.06 24.30 18.37 11.37 11.03 27.80 19.00 10.78
CHIPS (ours) 9.14 13.82 24.70 23.67 14.28 13.67 29.82 13.00 10.36
B16-400M
Random 6.57 11.27 23.00 13.80 10.35 10.49 25.75 17.75 4.68
TracIn 6.53 12.02 18.00 13.98 12.85 15.28 22.98 40.50 4.96
CHIPS (ours) 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
B16-$C$
Random 9.92 11.82 25.50 17.61 11.04 12.47 31.50 6.00 5.38
TracIn 9.12 12.82 23.30 14.47 14.05 15.37 41.14 8.50 5.94
CHIPS (ours) 10.10 12.37 24.20 13.61 13.56 14.42 43.15 9.75 6.22
L14-400M
Random 9.81 19.10 23.20 26.17 16.99 17.37 36.39 38.50 23.02
TracIn 10.51 21.05 23.30 23.10 16.13 17.55 36.92 31.00 11.04
CHIPS (ours) 8.74 25.74 22.20 23.01 15.06 17.07 35.04 38.50 17.37
L14-$C$
Random 13.06 21.40 26.20 25.68 19.81 19.23 30.74 20.25 7.31
TracIn 15.82 27.83 23.60 16.71 15.68 11.29 21.88 20.25 6.11
CHIPS (ours) 16.39 28.99 23.30 19.39 16.38 15.26 25.42 16.75 7.74
H14-$C$
Random 8.97 62.19 24.80 25.52 16.69 14.35 38.29 8.00 4.83
TracIn 13.05 64.69 23.80 31.44 20.53 18.83 37.51 22.25 6.53
CHIPS (ours) 8.89 54.16 23.20 27.00 18.78 15.67 34.21 19.25 7.24
Table 13: Full medical-domain results (Part 2/2) of the generalization experiment.
Model Cars Country211 FER Aircraft Food101 GTSRB
Imagenet
-A
Imagenet
-O
Imagenet
1k
Imagenet
v2
B32-400M
Random 66.66 16.73 30.54 25.44 78.84 38.45 27.85 45.80 63.49 55.78
TracIn 66.00 15.50 33.74 22.41 77.07 27.69 27.07 40.75 61.35 52.69
CHIPS (ours) 65.97 15.85 31.51 24.42 77.74 30.26 28.40 42.10 62.46 53.85
B32-$C$
Random 74.99 17.20 43.52 23.34 80.65 39.71 28.37 46.35 65.90 57.95
TracIn 73.96 15.58 35.82 23.46 78.49 37.19 29.13 40.85 63.71 55.60
CHIPS (ours) 73.37 16.05 45.75 24.15 79.41 35.33 29.73 41.55 64.16 56.40
B16-400M
Random 71.43 21.76 43.73 27.06 84.99 41.94 45.60 39.20 69.18 61.31
TracIn 70.63 19.09 37.11 23.91 81.94 31.37 40.48 31.20 63.62 55.34
CHIPS (ours) 68.40 21.47 40.69 25.32 82.60 31.02 42.36 32.50 66.88 56.22
B16-$C$
Random 81.12 22.35 48.75 30.42 86.97 51.00 48.59 40.85 71.06 64.12
TracIn 75.25 17.76 34.20 22.59 81.33 36.29 41.97 31.20 62.65 55.68
CHIPS (ours) 75.96 18.10 35.33 23.49 82.34 40.96 43.88 32.20 64.15 57.34
L14-400M
Random 82.93 30.34 41.36 38.97 89.35 48.79 65.73 29.35 75.23 68.72
TracIn 78.54 26.27 36.57 35.25 87.81 39.90 61.49 24.30 69.84 63.35
CHIPS (ours) 80.26 26.73 39.45 34.71 88.02 41.71 62.24 25.20 70.88 64.61
L14-$C$
Random 88.02 33.86 54.99 42.18 93.31 56.33 71.47 29.05 78.28 71.32
TracIn 83.62 29.36 46.68 41.19 90.38 48.76 66.41 24.65 73.51 66.43
CHIPS (ours) 84.65 29.78 47.94 40.98 91.25 46.98 68.07 25.15 74.15 67.55
H14-$C$
Random 89.64 37.46 52.49 51.43 93.87 58.34 74.92 29.45 79.75 73.69
TracIn 88.73 35.35 53.19 45.15 92.81 57.78 73.16 24.05 76.19 69.91
CHIPS (ours) 88.30 35.64 55.46 47.58 93.09 57.86 74.25 24.85 76.63 70.39
Table 14: Full general-domain results (Part 1/4) of the generalization experiment.
Model MNIST
Rendered
SST2
STL10 Sun397
Sun397
Official
VOC Caltech101 CIFAR10 CIFAR100
B32-400M
Random 39.99 54.81 96.11 63.75 64.66 66.37 84.21 91.19 67.92
TracIn 35.04 53.76 94.42 59.56 60.19 68.32 82.86 88.24 63.02
CHIPS (ours) 38.52 55.13 95.08 59.43 60.70 66.67 82.70 88.68 63.70
B32-$C$
Random 43.11 53.87 96.42 66.50 66.95 76.30 85.80 95.17 77.92
TracIn 38.90 52.83 95.40 63.29 64.03 75.01 86.06 94.27 76.24
CHIPS (ours) 36.16 51.62 95.71 63.95 64.63 75.73 84.62 94.01 75.47
B16-400M
Random 45.07 55.85 97.05 66.82 67.89 69.95 85.55 91.94 67.61
TracIn 39.09 52.94 95.24 57.33 59.55 68.56 82.17 82.60 54.60
CHIPS (ours) 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
B16-$C$
Random 59.17 54.75 98.35 68.43 69.29 78.17 84.54 96.09 79.97
TracIn 54.70 53.27 97.12 58.73 60.19 69.29 82.28 92.97 72.27
CHIPS (ours) 55.66 54.42 97.61 61.01 62.31 70.19 82.68 93.50 73.19
L14-400M
Random 52.89 64.52 99.20 71.34 72.54 74.47 85.98 95.97 76.67
TracIn 58.40 58.26 97.34 62.08 63.28 62.04 82.28 89.52 70.96
CHIPS (ours) 59.23 62.93 97.74 61.34 62.34 62.95 82.83 91.16 70.80
L14-$C$
Random 60.93 68.26 99.25 72.03 73.77 80.58 88.15 97.60 84.84
TracIn 57.61 56.73 98.61 66.52 67.98 74.63 84.45 95.20 79.35
CHIPS (ours) 64.90 64.80 98.84 67.41 68.64 75.45 84.42 95.46 78.86
H14-$C$
Random 70.45 70.68 99.49 73.01 75.34 73.45 87.54 98.10 86.61
TracIn 33.99 66.89 99.19 68.30 69.37 67.22 83.43 97.15 84.05
CHIPS (ours) 20.47 63.87 99.20 68.98 70.47 69.35 85.18 97.09 84.10
Table 15: Full general-domain results (Part 2/4) of the generalization experiment.
Model
CLEVR
Closest
CLEVR
Count
DMLAB DTD Eurosat Flowers KITTI
B32-400M
Random 21.01 23.33 19.06 50.32 50.06 70.06 32.63
TracIn 21.85 22.33 16.89 47.55 47.96 66.60 24.61
CHIPS (ours) 22.40 23.06 15.15 48.30 45.43 69.15 32.07
B32-$C$
Random 23.78 21.09 12.20 55.64 49.52 69.07 15.33
TracIn 22.59 19.37 12.34 53.03 45.76 63.77 15.61
CHIPS (ours) 22.55 20.12 13.17 54.52 43.00 66.11 19.69
B16-400M
Random 22.57 29.39 12.28 53.56 53.31 72.68 31.08
TracIn 21.39 22.96 14.83 45.80 40.17 66.08 17.02
CHIPS (ours) 22.59 21.31 14.92 48.86 40.27 66.29 19.83
B16-$C$
Random 22.55 28.51 21.39 61.65 52.61 74.94 27.00
TracIn 22.69 25.08 19.25 51.60 53.44 62.06 22.22
CHIPS (ours) 22.72 26.79 19.38 52.87 55.43 65.07 22.08
L14-400M
Random 21.07 34.80 20.12 59.84 62.15 77.90 28.41
TracIn 22.93 27.41 19.26 51.97 50.56 73.78 30.38
CHIPS (ours) 23.67 27.62 17.88 53.72 50.19 75.05 29.54
L14-$C$
Random 22.54 28.61 16.64 66.38 66.76 81.53 24.33
TracIn 22.54 27.45 20.95 59.89 56.39 78.09 25.04
CHIPS (ours) 22.54 28.74 18.50 60.37 53.33 78.60 32.77
H14-$C$
Random 10.25 20.78 14.86 68.99 69.37 83.72 27.14
TracIn 21.85 23.17 13.62 63.24 62.37 81.18 22.93
CHIPS (ours) 20.85 21.61 13.02 66.17 61.69 81.43 25.32
Table 16: Full general-domain results (Part 3/4) of the generalization experiment.
Model Pets RESISC45
Smallnorb
Azimuth
Smallnorb
Elevation
SVHN
B32-400M
Random 85.75 57.11 5.35 12.03 23.40
TracIn 84.79 55.87 5.74 11.45 27.67
CHIPS (ours) 86.67 56.70 5.24 12.04 25.37
B32-$C$
Random 88.74 59.37 5.70 12.20 18.77
TracIn 87.84 56.25 5.60 11.65 22.40
CHIPS (ours) 88.47 58.83 6.09 12.47 21.03
B16-400M
Random 90.52 64.08 5.93 10.88 18.41
TracIn 88.23 60.14 5.47 11.37 24.96
CHIPS (ours) 86.97 58.44 5.55 11.82 25.23
B16-$C$
Random 91.09 66.79 4.95 10.77 37.11
TracIn 87.57 59.35 5.33 11.01 36.07
CHIPS (ours) 87.82 60.21 5.28 10.77 36.69
L14-400M
Random 92.89 68.68 5.51 11.04 22.41
TracIn 90.30 63.06 5.93 11.94 27.65
CHIPS (ours) 91.41 63.32 5.85 12.17 27.55
L14-$C$
Random 94.06 75.13 4.85 11.13 46.73
TracIn 90.81 66.89 5.40 11.11 49.62
CHIPS (ours) 92.37 66.44 5.21 11.56 50.27
H14-$C$
Random 95.53 72.56 6.30 12.30 44.79
TracIn 94.19 68.65 5.82 12.40 50.89
CHIPS (ours) 94.49 68.44 6.44 12.21 50.99
Table 17: Full general-domain results (Part 4/4) of the generalization experiment.
H.3 Ablation Experiment
The full results of the main experiment are detailed in Tab. 18 and Tab. 19 for the medical domain, and in Tab. 20 through Tab. 23 for the general domain.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
r=10%
Alignment-only 59.51 57.64 39.41 12.76 88.69 13.17 23.62 3.71
Alignment-Margin 59.56 57.65 40.21 11.69 87.51 13.94 23.56 3.26
CHIPS (ours) 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
r=20%
Alignment-only 72.41 68.04 46.98 30.54 55.27 21.32 20.56 12.24
Alignment-Margin 72.02 68.32 46.87 30.10 58.24 21.65 21.34 12.34
CHIPS (ours) 72.50 68.67 48.07 30.36 53.58 23.50 22.49 13.97
r=30%
Alignment-only 71.30 72.76 43.79 32.14 60.26 22.71 21.88 11.77
Alignment-Margin 71.24 70.43 48.16 32.42 61.79 25.12 23.44 13.19
CHIPS (ours) 73.42 73.66 49.88 36.34 67.90 25.15 25.02 14.39
Table 18: Full medical-domain results (Part 1/2) of the ablation experiment.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
r=10%
Alignment-only 21.49 12.07 20.20 11.65 10.08 10.81 27.28 27.00 4.86
Alignment-Margin 21.02 12.17 19.60 12.10 10.00 10.28 26.70 29.25 4.99
CHIPS (ours) 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
r=20%
Alignment-only 9.57 11.22 20.80 14.10 13.86 15.24 24.57 32.00 5.99
Alignment-Margin 9.07 11.12 20.80 14.53 14.17 15.04 24.55 35.75 5.39
CHIPS (ours) 8.96 11.87 21.40 13.50 13.80 14.87 23.20 38.50 5.92
r=30%
Alignment-only 8.28 11.22 21.80 13.62 12.99 14.87 21.56 30.75 7.46
Alignment-Margin 10.49 11.32 20.50 14.35 13.38 12.69 24.08 30.50 6.40
CHIPS (ours) 10.65 11.37 21.30 13.13 14.28 15.33 24.43 30.25 7.69
Table 19: Full medical-domain results (Part 2/2) of the ablation experiment.
Model Cars Country211 FER Aircraft Food101 GSTRB
Imagenet
-A
Imagenet
-O
Imagenet
1k
Imagenet
v2
r=10%
Alignment-only 69.99 19.50 40.85 25.59 82.55 31.56 42.47 32.95 64.84 56.63
Alignment-Margin 70.12 21.67 40.81 25.92 82.60 31.62 42.17 32.45 65.01 56.78
CHIPS (ours) 68.40 21.47 40.69 25.32 82.60 31.02 42.36 32.50 66.88 56.22
r=20%
Alignment-only 68.97 18.80 39.19 24.60 80.06 31.50 40.11 31.60 62.95 55.05
Alignment-Margin 68.95 19.59 39.23 24.51 80.03 31.92 40.15 31.60 62.89 54.99
CHIPS (ours) 69.32 19.54 39.40 24.36 80.05 31.06 40.33 32.10 62.97 55.01
r=30%
Alignment-only 68.86 18.41 38.35 24.78 79.45 30.78 39.45 31.60 62.63 54.52
Alignment-Margin 68.74 18.48 42.28 24.18 79.93 30.43 39.83 31.40 62.56 54.74
CHIPS (ours) 68.30 18.39 37.89 23.46 79.27 30.36 38.96 31.40 62.33 54.45
Table 20: Full general-domain results (Part 1/4) of the ablation experiment.
Model MNIST
Rendered
SST2
STL10 Sun397
Sun397
Official
VOC Caltech101 CIFAR10 CIFAR100
r=10%
Alignment-only 45.04 56.78 95.42 58.44 60.47 66.18 83.27 82.55 55.19
Alignment-Margin 44.47 57.44 95.34 58.34 60.40 66.66 82.93 82.57 55.34
CHIPS (ours) 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
r=20%
Alignment-only 43.42 50.74 95.19 56.52 58.73 66.31 81.38 84.85 57.73
Alignment-Margin 42.65 50.36 94.99 56.47 58.76 65.93 81.10 85.76 59.11
CHIPS (ours) 42.91 50.63 95.42 55.94 58.51 65.89 81.38 85.72 59.00
r=30%
Alignment-only 38.92 50.14 95.21 56.16 58.28 65.77 80.12 85.44 57.97
Alignment-Margin 35.97 50.03 95.17 55.95 58.43 62.85 80.30 84.61 56.85
CHIPS (ours) 34.25 49.92 95.36 55.69 57.91 64.36 80.79 84.61 57.64
Table 21: Full general-domain results (Part 2/4) of the ablation experiment.
Model
CLEVR
Closest
CLEVR
Count
DMLAB DTD Eurosat Flowers KITTI
r=10%
Alignment-only 22.62 21.30 13.90 48.67 38.98 68.11 23.49
Alignment-Margin 22.58 21.41 14.14 49.36 39.11 68.01 23.21
CHIPS (ours) 22.59 21.31 14.92 48.86 40.27 66.29 19.83
r=20%
Alignment-only 21.27 24.39 15.90 44.79 41.89 66.22 17.86
Alignment-Margin 21.00 25.33 15.90 44.26 42.15 66.30 17.44
CHIPS (ours) 21.52 25.25 16.21 44.47 42.89 66.53 16.46
r=30%
Alignment-only 21.24 21.56 15.40 45.21 42.76 65.03 15.61
Alignment-Margin 22.24 20.59 16.86 44.15 42.06 65.67 17.72
CHIPS (ours) 21.60 20.78 15.76 44.20 42.19 65.30 14.77
Table 22: Full general-domain results (Part 3/4) of the ablation experiment.
Model Pets RESISC45
Smallnorb
Azimuth
Smallnorb
Elevation
SVHN
r=10%
Alignment-only 87.46 60.33 5.51 11.34 26.19
Alignment-Margin 87.35 60.21 5.47 11.32 26.04
CHIPS (ours) 86.97 58.44 5.55 11.82 25.23
r=20%
Alignment-only 87.35 59.48 5.62 11.06 24.34
Alignment-Margin 87.35 59.49 5.32 10.74 23.96
CHIPS (ours) 87.24 59.68 5.49 11.02 24.63
r=30%
Alignment-only 86.70 59.60 5.53 10.79 23.94
Alignment-Margin 87.00 59.68 5.82 10.99 23.79
CHIPS (ours) 87.08 60.00 5.84 11.18 22.45
Table 23: Full general-domain results (Part 4/4) of the ablation experiment.
H.4 Analysis Experiment
  • Hyperparameter Analysis: The analysis of evaluation set size, mixing hyperparameter $\alpha$, and balance hyperparameter $\beta$ is presented in Tab. 24 through Tab. 29.
  • End-point Subspace: The analysis of the end-point subspace $\vartheta$ is shown in Tab. 30 and Tab. 31.
  • JL Random Projection: The analysis of JL random projection is provided in Tab. 32 and Tab. 33.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
Evaluation Set Size
50 67.33 57.66 39.68 18.10 87.22 12.69 21.51 2.79
100 57.57 58.45 39.39 14.90 82.92 13.02 24.76 2.50
150 54.62 56.38 40.61 11.38 85.64 13.07 24.06 4.61
200 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
250 68.07 66.49 40.55 22.20 87.99 12.63 23.91 2.14
$\alpha$
0.2 52.04 57.57 40.53 13.70 79.23 13.61 24.55 4.28
0.4 51.45 57.18 41.20 13.51 74.95 14.03 24.90 3.28
0.6 56.65 55.74 40.21 14.39 89.81 13.85 25.17 2.90
0.8 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
1.0 51.01 55.88 40.67 14.77 83.71 13.51 23.79 3.29
$\beta$
0 71.38 59.27 44.43 27.03 43.02 25.05 18.36 3.13
0.25 54.72 56.48 39.89 11.75 90.43 12.74 22.42 3.08
0.50 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
0.75 52.53 57.27 38.93 13.83 86.47 13.79 23.21 4.25
1.0 71.53 59.11 41.89 26.40 54.35 23.21 15.76 2.71
Table 24: Full medical-domain results (Part 1/2) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
Evaluation Set Size
50 16.07 12.07 21.50 8.66 8.70 10.04 26.56 35.25 5.20
100 21.20 12.17 20.70 9.23 8.87 9.49 26.24 32.75 5.87
150 28.48 12.32 21.00 11.42 10.32 10.59 30.21 35.00 4.92
200 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
250 16.49 12.07 19.50 12.48 10.13 9.88 24.38 30.25 5.02
$\alpha$
0.2 30.17 12.12 20.30 10.93 9.54 10.30 28.23 30.00 4.92
0.4 29.49 12.17 20.70 10.98 10.08 10.82 28.08 29.25 4.82
0.6 28.46 12.32 20.70 11.76 10.66 11.74 29.33 33.00 4.95
0.8 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
1.0 28.75 12.07 20.30 11.49 10.30 11.39 29.04 30.50 5.04
$\beta$
0 7.56 11.32 19.90 10.15 11.19 12.91 24.43 37.00 5.80
0.25 19.36 12.07 20.60 11.63 10.02 10.11 26.38 35.50 5.18
0.50 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
0.75 22.38 11.82 19.90 11.30 10.70 11.05 29.28 32.75 5.45
1.0 6.39 10.97 19.50 10.09 11.51 12.71 24.32 38.75 5.68
Table 25: Full medical-domain results (Part 2/2) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model Cars Country211 FER Aircraft Food101 GSTRB
Imagenet
-A
Imagenet
-O
Imagenet
1k
Imagenet
v2
Evaluation Set Size
50 70.15 19.38 40.58 25.65 82.56 31.46 42.68 33.05 65.04 57.01
100 69.79 19.40 41.70 25.26 82.52 33.02 42.25 32.10 65.15 57.00
150 69.48 19.48 41.07 24.75 82.70 32.54 42.09 33.40 64.97 56.96
200 68.40 21.47 40.69 25.32 82.60 31.02 42.36 32.50 66.88 56.22
250 67.38 19.41 40.35 25.08 82.91 30.02 42.45 32.25 64.91 56.80
$\alpha$
0.2 69.59 19.26 41.71 24.99 82.4 32.38 42.72 32.35 64.97 56.81
0.4 69.82 19.34 40.97 25.26 82.65 32.07 42.71 32.40 65.12 57.15
0.6 68.40 21.47 40.69 25.32 82.60 31.02 42.36 32.50 66.88 56.22
0.8 69.88 19.40 40.92 25.77 82.56 31.95 42.85 31.70 65.08 57.13
1.0 70.07 19.50 41.31 25.56 82.63 31.50 43.00 32.15 65.12 57.08
$\beta$
0 71.00 19.63 42.77 25.77 82.05 32.64 42.29 31.70 64.59 56.71
0.25 69.98 19.42 40.14 25.26 82.51 31.57 41.99 32.30 64.99 56.68
0.50 68.40 21.47 40.69 25.32 82.60 31.02 42.36 32.50 66.88 56.22
0.75 70.54 19.26 41.39 25.26 82.66 33.66 42.52 32.10 65.11 57.02
1.0 70.39 19.41 41.84 26.76 81.91 32.07 41.99 32.25 64.63 56.26
Table 26: Full general-domain results (Part 1/4) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model MNIST
Rendered
SST2
STL10 Sun397
Sun397
Official
VOC Caltech101 CIFAR10 CIFAR100
Evaluation Set Size
50 47.03 57.77 95.29 58.61 60.61 66.67 82.74 82.78 54.91
100 46.74 56.84 95.40 58.71 60.62 65.63 83.47 82.48 54.73
150 43.58 57.00 95.31 59.10 60.79 64.98 83.57 82.77 55.25
200 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
250 42.26 58.26 95.09 57.58 60.68 63.62 83.25 80.11 55.04
$\alpha$
0.2 45.55 57.44 95.31 58.91 60.49 66.91 83.06 82.61 55.31
0.4 45.44 57.83 95.40 58.78 60.65 67.05 83.11 83.08 55.68
0.6 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
0.8 45.82 57.33 95.39 58.92 60.55 66.19 83.27 82.96 55.15
1.0 47.11 57.66 95.51 58.73 60.53 66.57 83.27 82.95 55.21
$\beta$
0 42.85 52.33 95.51 57.74 60.47 66.73 82.65 85.81 58.97
0.25 44.89 58.26 94.31 58.89 60.89 66.29 82.97 82.37 54.13
0.50 38.82 57.84 94.91 58.00 57.88 68.66 82.48 80.76 55.33
0.75 42.92 58.37 95.34 58.86 60.79 66.25 83.58 84.02 56.88
1.0 40.30 51.89 95.14 57.72 60.44 66.75 82.51 85.63 59.00
Table 27: Full general-domain results (Part 2/4) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model
CLEVR
Closest
CLEVR
Count
DMLAB DTD Eurosat Flowers KITTI
Evaluation Set Size
50 22.65 21.22 14.08 48.24 39.11 68.11 23.21
100 22.65 21.45 14.08 48.83 38.69 68.04 22.22
150 22.60 22.14 14.63 48.35 39.52 67.47 23.07
200 22.59 21.31 14.92 48.86 40.27 66.29 19.83
250 22.44 21.19 13.76 48.35 37.52 67.05 25.32
$\alpha$
0.2 22.59 21.78 14.44 47.93 38.48 68.21 22.22
0.4 22.59 21.87 14.42 48.51 38.28 68.14 22.78
0.6 22.59 21.31 14.92 48.86 40.27 66.29 19.83
0.8 22.61 21.41 14.03 48.40 38.52 67.99 21.94
1.0 22.63 21.91 14.18 48.62 38.28 67.99 21.10
$\beta$
0 22.10 26.25 15.35 46.91 42.26 68.35 20.53
0.25 22.66 21.09 13.89 48.14 39.11 68.48 23.07
0.50 22.59 21.31 14.92 48.86 40.27 66.29 19.83
0.75 22.55 21.87 13.98 47.93 37.81 67.75 23.21
1.0 22.06 26.15 15.76 47.23 42.78 68.24 20.53
Table 28: Full general-domain results (Part 3/4) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model Pets RESISC45
Smallnorb
Azimuth
Smallnorb
Elevation
SVHN
Evaluation Set Size
50 87.63 60.05 5.44 11.27 25.64
100 88.06 60.27 5.51 11.51 26.43
150 88.01 60.24 5.56 11.77 26.67
200 86.97 58.44 5.55 11.82 25.23
250 87.49 60.38 5.42 11.32 26.01
$\alpha$
0.2 87.76 60.56 5.42 11.41 25.56
0.4 87.46 60.46 5.79 11.49 25.97
0.6 86.97 58.44 5.55 11.82 25.23
0.8 87.74 60.62 5.28 11.49 26.41
1.0 87.90 60.43 5.33 11.28 26.49
$\beta$
0 87.93 62.27 5.77 10.63 25.60
0.25 87.74 60.14 5.57 11.54 25.85
0.50 86.97 58.44 5.55 11.82 25.23
0.75 87.44 59.79 5.72 11.55 26.83
1.0 87.27 61.63 5.73 11.06 25.86
Table 29: Full general-domain results (Part 4/4) of the analysis on evaluation set size, $\alpha$, and $\beta$.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
All
Dot 71.67 57.20 39.89 23.51 72.64 18.34 16.87 5.50
TracIn 65.40 56.75 40.03 26.34 90.73 12.84 16.69 3.69
TRAK 43.23 59.23 39.57 12.88 77.16 14.74 25.93 3.60
CHIPS (ours) 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
Logit-only
Dot 62.45 59.27 31.28 25.96 13.48 63.18 20.90 9.60
TracIn 58.24 61.00 33.89 16.09 8.30 71.48 19.38 10.15
TRAK 64.29 62.24 35.12 26.02 14.73 59.51 17.36 11.74
CHIPS (ours) 59.94 61.30 35.36 20.93 13.72 70.29 17.89 21.50
Visual-only
Dot 61.96 56.15 37.71 23.00 13.33 47.26 20.05 6.74
TracIn 59.11 53.51 34.99 18.16 28.44 31.75 18.15 13.78
TRAK 71.23 59.48 44.83 32.75 23.59 24.27 15.02 15.37
CHIPS (ours) 64.01 57.61 41.25 42.61 33.29 53.67 20.93 5.48
Text-only
Dot 71.29 61.81 40.77 24.07 40.91 12.56 17.10 20.57
TracIn 61.86 60.90 39.23 25.71 77.58 13.00 18.65 11.51
TRAK 51.61 57.08 39.55 21.36 34.21 46.31 20.61 11.22
CHIPS (ours) 69.76 60.02 46.83 35.80 25.76 55.13 18.65 12.80
Table 30: Full medical-domain results (Part 1/2) of the analysis on end-point subspace.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
All
Dot 6.17 12.17 20.60 9.48 10.36 11.19 25.42 33.75 4.16
TracIn 6.53 12.02 18.00 13.98 12.85 15.28 22.98 40.50 4.96
TRAK 25.84 12.17 20.90 10.74 10.43 11.22 28.68 26.75 4.46
CHIPS (ours) 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
Logit-only
Dot 9.63 11.72 24.10 9.22 9.10 7.31 13.89 21.50 6.13
TracIn 10.82 11.77 22.40 10.54 8.82 7.50 15.08 22.25 5.21
TRAK 6.84 11.92 22.80 9.20 7.80 7.34 15.10 23.25 5.53
CHIPS (ours) 9.49 12.02 22.50 9.42 8.68 7.68 16.18 19.00 5.16
Visual-only
Dot 3.66 11.62 21.30 11.04 9.30 10.72 24.21 22.75 4.72
TracIn 6.45 10.97 19.80 11.80 12.04 13.90 24.25 26.00 4.95
TRAK 6.20 11.77 18.90 11.51 9.90 10.42 17.19 40.00 5.56
CHIPS (ours) 19.32 12.02 19.60 10.91 10.69 12.22 23.69 26.50 5.31
Text-only
Dot 11.54 11.97 22.50 11.72 11.81 12.30 23.79 33.00 4.55
TracIn 10.52 12.07 19.80 10.27 12.34 13.11 23.97 40.25 4.69
TRAK 18.09 12.27 22.30 12.52 11.82 12.82 25.97 27.25 6.31
CHIPS (ours) 19.51 11.42 19.70 10.09 9.31 9.67 17.66 39.25 5.09
Table 31: Full medical-domain results (Part 2/2) of the analysis on end-point subspace.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
CountSketch
2k 65.13 59.60 48.05 21.43 35.20 32.39 16.69 4.16
4k 68.35 63.66 40.51 25.96 88.06 12.38 21.43 6.24
8k 63.37 57.97 39.76 17.03 73.36 14.87 22.45 2.14
16k 62.10 55.73 40.00 21.50 79.72 23.67 22.27 3.38
Sparse
2k 66.86 61.11 40.91 25.27 18.14 68.09 17.45 5.81
4k 68.68 60.15 44.19 23.76 23.67 39.46 10.38 5.01
8k 65.29 65.76 43.33 27.03 41.30 12.49 14.53 4.12
16k 68.51 62.10 43.89 35.89 43.41 56.96 24.26 8.14
SRHT
2k 32.69 60.39 47.09 33.19 68.64 39.88 18.91 6.90
4k 69.35 58.69 39.04 14.39 88.25 15.72 18.65 8.84
8k 41.47 56.59 39.31 27.66 88.67 20.62 14.21 1.82
16k 65.89 58.36 40.35 41.04 32.65 44.25 12.72 3.77
Table 32: Full medical-domain results (Part 1/2) of the analysis on JL random projection.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
CountSketch
2k 10.10 11.32 17.80 14.33 15.17 16.28 21.18 28.00 5.63
4k 16.63 11.87 18.20 12.03 9.64 10.71 23.90 30.25 5.29
8k 21.02 12.77 17.40 14.20 14.00 13.61 27.45 30.50 5.88
16k 20.71 12.37 21.40 14.04 14.62 13.73 25.96 36.00 8.09
Sparse
2k 11.39 11.12 23.90 14.60 13.47 13.90 26.13 35.75 6.81
4k 10.47 11.72 20.60 10.70 11.31 11.35 24.35 30.00 6.68
8k 7.93 12.17 15.10 11.33 12.20 15.33 23.16 24.50 5.33
16k 7.06 11.42 20.70 12.30 12.07 12.35 28.83 31.75 5.60
SRHT
2k 13.51 11.57 22.50 10.75 10.47 10.39 23.23 27.00 5.27
4k 9.45 12.02 17.10 9.60 10.72 11.03 31.63 36.00 6.47
8k 22.18 11.87 17.40 13.21 13.21 13.23 28.70 27.00 5.91
16k 7.81 11.22 18.30 12.30 12.48 11.63 24.75 26.25 6.55
Table 33: Full medical-domain results (Part 2/2) of the analysis on JL random projection.
Appendix I Additional Results
We provide additional experiments on MedTrinity [56] in Tab. 34 and Tab. 35.
Model Diabetic PCAM LC25000 Pollen Amyloid CAA Amyloid Diffuse BloodMNIST ChestMNIST
B32-400M
Random 10% 63.65 50.20 32.51 72.91 97.26 12.82 14.88 6.86
Random 50% 34.53 52.70 25.68 72.91 97.76 12.50 9.97 28.41
B32-$C$
Random 10% 3.45 63.08 8.80 11.00 1.84 87.94 29.49 2.75
Random 50% 3.11 56.25 23.89 51.85 2.86 87.94 21.60 17.87
B16-400M
Random 10% 10.74 56.77 39.28 7.23 88.27 59.82 14.44 9.08
Random 50% 17.08 66.73 32.91 8.30 96.46 14.03 7.28 1.24
B16-$C$
Random 10% 46.68 50.49 6.40 21.06 60.28 87.95 6.58 1.56
Random 50% 22.16 50.83 17.95 61.28 62.36 87.95 8.21 1.07
Table 34: Medical-domain results (Part 1/2) of MedTrinity [56] dataset.
Model ChestXray14 Derma Oct OrganA OrganC OrganS Path Retina Tissue
B32-400M
Random 10% 6.05 13.97 25.00 10.77 6.86 5.30 34.50 9.75 5.31
Random 50% 11.29 11.32 22.60 10.68 6.69 6.30 35.72 8.50 5.02
B32-$C$
Random 10% 4.42 9.63 25.00 16.22 8.98 6.24 31.91 14.25 9.18
Random 50% 7.56 11.07 25.00 21.59 14.56 14.72 46.03 16.00 8.01
B16-400M
Random 10% 4.78 10.62 12.40 13.98 8.57 7.73 27.73 12.75 9.46
Random 50% 2.86 10.62 16.20 13.17 11.93 10.32 30.28 7.50 9.70
B16-$C$
Random 10% 9.89 10.62 25.90 13.04 12.71 13.33 35.65 5.00 6.15
Random 50% 6.45 18.50 25.40 19.92 14.79 15.54 42.45 5.75 4.92
Table 35: Medical-domain results (Part 2/2) of MedTrinity [56] dataset.